The Not So Short Introduction to L^{A}TEX 2ε

Or L^{A}TEX2e in 69 minutes

by Tobias Oetiker

Hubert Partl, Irene Hyna and Elisabeth Schlegl Version 2.2-Beta, 25th January, 1996 \underline{ii}

This document is “public domain”. It may be printed and distributed free of charge in its original form (including the list of authors). If it is changed or if parts of it are used within another document, then the author list must include all the original authors AND that author (those authors) who has (have) made the changes. When this document is used commercially, the GNU copyleft applies.

Thank you!

Much of the material used in this introduction comes from an Austrian introduction to L^{A}TEX 2.09 written in German by:

Hubert Partl <partl@mail.boku.ac.at>

Zentraler Informatikdienst der Universität für Bodenkultur Wien

Irene Hyna <Irene.Hyna@bmwf.ac.at>

Bundesministerium für Wissenschaft und Forschung Wien

Elisabeth Schlegl <no email>

in Graz

If you are interested in the German document you can find a version updated for L^{A}TEX 2ε by Jörg Knappen at CTAN:/tex-archive/info/lkurz

While preparing this document I asked for reviewers on comp.text.tex. I got a lot of response. The following individuals helped with corrections, suggestions and material to improve this paper. They put in a big effort to help me get this document into its present shape. I would like to sincerely thank all of them. Naturally, all the mistakes you’ll find in this book are mine. If you ever find a word which is spelled correctly, it must have been one of the people below dropping me a line.

Rosemary Bailey, David Carlisle, Chris McCormack, David Dureisseix, Elliot, Robin Fairbairns, Alexandre Guimond, Cyril Goutte, Neil Hammond, Rasmus Borup Hansen, Martien Hulsen, Eric Jacoboni, Alan Jeffrey, Byron Jones, David Jones, Andrzej Kawalec, Christian Kern, Jörg Knappen, Maik Lehradt, Claus Malten, Hubert Partl, John Refling, Mike Ressler, Brian Ripley, Young U. Ryu, Chris Rowley, Craig Schlenter, and Josef Tkadlec.

Preface

L^{A}TEX[1] is a typesetting system which is most suited to producing scientific and mathematical documents of high typographical quality. The system is also suitable for producing all sorts of other documents, from simple letters to complete books. L^{A}TEX uses TEX[2] as its formatting engine.

This short introduction describes L^{A}TEX 2ε and should be sufficient for most applications of L^{A}TEX. For a complete description of the L^{A}TEX system refer to [1, 3].

L^{A}TEX is available for most computers from the IBM PC upwards. On many university computer networks the system is already installed, ready to operate. Information on how to access the local L^{A}TEX installation should be provided in the Local Guide [4]. If you have problems getting started, ask the person who gave you this booklet. The scope of this document is not to tell you how to install and set up a L^{A}TEX system, but to teach you how to write your documents so that they can be processed by L^{A}TEX.

This Introduction is split into 4 chapters:

Chapter 1 tells you about the basic structure of L^{A}TEX 2ε documents. You will also learn a bit about the history of L^{A}TEX. After reading this chapter you should have a rough picture of L^{A}TEX. The picture will only be a framework, but it will enable you to integrate the information provided in the other chapters into the big picture.

Chapter 2 goes into the details of typesetting your documents. It explains most of the essential L^{A}TEX commands and environments. After reading this chapter you will be able to write your first documents.

Chapter 3 explains how to typeset formulae with L^{A}TEX. Again a lot of examples help you to understand how to use one of L^{A}TEX main strengths. At the end of this chapter you find tables, listing all the mathematical symbols available in L^{A}TEX.

Chapter 4 adds in some bits and bobs about L^{A}TEX which are not essential, but very handy. Among other things you will learn how to include eps graphics into your documents or how to add an index to your publication.

It is important to read the chapters in sequential order. The book is not that big after all. Make sure to carefully read the examples, because a great part of the information is contained in the various examples you will find all through out the book.

If you need to get hold of any L^{A}TEX related material, have a look in one of the CTAN ftp archives. For the US it is at ftp.shsu.edu, for Germany it is ftp.dante.de and for the UK it is ftp.tex.ac.uk. If you are not in one of these countries, choose the archive closest to you.

If you have ideas for something to be added, removed or altered in this document, please let me know. I am especially interested in feedback from L^{A}TEX novices about which bits of this intro are easy to understand and which could be explained better.

Tobias Oetiker <oetiker@ee.ethz.ch>

Department of Electrical Engineering, Swiss Federal Institute of Technology

The current version of this document will be available on

CTAN:/tex-archive/info/lshort

Contents

Thank you! iii

Preface v

1 Things You Need to Know 1

1.1 The Name of the Game . . . . . . . . . . . . . . . . . . . . . 1 1.1.1 TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.1.2 L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.2 Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2.1 Author, Book Designer, and Typesetter . . . . . . . . 2 1.2.2 Layout Design . . . . . . . . . . . . . . . . . . . . . . 2 1.2.3 Advantages and Disadvantages . . . . . . . . . . . . . 3

1.3 L^{A}TEX Input Files . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.1 Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.2 Special Characters . . . . . . . . . . . . . . . . . . . . 4 1.3.3 L^{A}TEX Commands . . . . . . . . . . . . . . . . . . . . 5 1.3.4 Comments . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.4 Input File Structure . . . . . . . . . . . . . . . . . . . . . . . 6 1.5 The Layout of the Document . . . . . . . . . . . . . . . . . . 7

1.5.1 Document Classes . . . . . . . . . . . . . . . . . . . . 7 1.5.2 Packages . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.5.3 Page Styles . . . . . . . . . . . . . . . . . . . . . . . . 11

1.6 Big Projects . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2 Typesetting Text 13

2.1 Linebreaking and Pagebreaking . . . . . . . . . . . . . . . . . 13 2.1.1 Justified Paragraphs . . . . . . . . . . . . . . . . . . . 13 2.1.2 Hyphenation . . . . . . . . . . . . . . . . . . . . . . . 14

2.2 Special Characters and Symbols . . . . . . . . . . . . . . . . . 15 2.2.1 Quotation Marks . . . . . . . . . . . . . . . . . . . . . 15 2.2.2 Dashes and Hyphens . . . . . . . . . . . . . . . . . . . 15 2.2.3 Ellipsis ( …) . . . . . . . . . . . . . . . . . . . . . . . 16 2.2.4 Ligatures . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.2.5 Accents and Special Characters . . . . . . . . . . . . . 16 2.3 International Language Support . . . . . . . . . . . . . . . . . 17 2.4 The Space between Words . . . . . . . . . . . . . . . . . . . . 18 2.5 Titles, Chapters, and Sections . . . . . . . . . . . . . . . . . . 18 2.6 Cross References . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7 Footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.8 Emphasized Words . . . . . . . . . . . . . . . . . . . . . . . . 21 2.9 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

2.9.1 Itemise, Enumerate, and Description . . . . . . . . . . 21 2.9.2 Flushleft, Flushright, and Center . . . . . . . . . . . . 22 2.9.3 Quote, Quotation, and Verse . . . . . . . . . . . . . . 22 2.9.4 Printing Verbatim . . . . . . . . . . . . . . . . . . . . 23 2.9.5 Tabular . . . . . . . . . . . . . . . . . . . . . . . . . . 24

2.10 Floating Bodies . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.11 Adding New Commands and Environments . . . . . . . . . . 28

2.11.1 New Commands . . . . . . . . . . . . . . . . . . . . . 28 2.11.2 New Environments . . . . . . . . . . . . . . . . . . . . 29

3 Typesetting Mathematical Formulae 31 3.1 General . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.2 Grouping in Math Mode . . . . . . . . . . . . . . . . . . . . . 33 3.3 Building Blocks of a Mathematical Formula . . . . . . . . . . 33 3.4 Math Spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.5 Vertically Aligned Material . . . . . . . . . . . . . . . . . . . 37 3.6 Math Font Size . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.7 Describing Variables . . . . . . . . . . . . . . . . . . . . . . . 39 3.8 Theorems, Laws, . . . . . . . . . . . . . . . . . . . . . . . . . . 40 3.9 Bold symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 3.10 List of Mathematical Symbols . . . . . . . . . . . . . . . . . . 42

4 Specialities 49 4.1 Fonts and Sizes . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.2 Spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

4.2.1 Line Spacing . . . . . . . . . . . . . . . . . . . . . . . 51 4.2.2 Paragraph Formating . . . . . . . . . . . . . . . . . . 51 4.2.3 Horizontal Space . . . . . . . . . . . . . . . . . . . . . 52 4.2.4 Vertical Space . . . . . . . . . . . . . . . . . . . . . . 53

4.3 Page Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.4 Bibliography . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.5 Indexing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 4.6 Including EPS Graphics . . . . . . . . . . . . . . . . . . . . . 57

List of Figures

1.1 A Minimal L^{A}TEX File . . . . . . . . . . . . . . . . . . . . . . 6 1.2 Example of a Realistic Journal Article . . . . . . . . . . . . . 7 4.1 Page Layout Parameters . . . . . . . . . . . . . . . . . . . . . 54

List of Tables

1.1 Document Classes . . . . . . . . . . . . . . . . . . . . . . . . 8 1.2 Document Class Options . . . . . . . . . . . . . . . . . . . . . 8 1.3 Some of the Packages Distributed with L^{A}TEX . . . . . . . . . 10 1.4 The Predefined Page Styles of L^{A}TEX . . . . . . . . . . . . . . 11

2.1 Accents and Special Characters . . . . . . . . . . . . . . . . . 17 2.2 Float Placing Permissions . . . . . . . . . . . . . . . . . . . . 26

3.1 Math Mode Accents . . . . . . . . . . . . . . . . . . . . . . . 42 3.2 Lowercase Greek Letters . . . . . . . . . . . . . . . . . . . . . 42 3.3 Uppercase Greek Letters . . . . . . . . . . . . . . . . . . . . . 42 3.4 Binary Relations . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.5 Binary Operators . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.6 BIG Operators . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.7 Arrows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.8 Delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.9 Large Delimiters . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.10 Miscellaneous Symbols . . . . . . . . . . . . . . . . . . . . . . 45 3.11 Non-Mathematical Symbols . . . . . . . . . . . . . . . . . . . 45 3.12 AMS Delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.13 AMS Greek and Hebrew . . . . . . . . . . . . . . . . . . . . . 45 3.14 AMS Binary Relations . . . . . . . . . . . . . . . . . . . . . . 46 3.15 AMS Arrows . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.16 AMS Negated Binary Relations and Arrows . . . . . . . . . . 47 3.17 AMS Binary Operators . . . . . . . . . . . . . . . . . . . . . 47 3.18 AMS Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . 48 3.19 Math Alphabets . . . . . . . . . . . . . . . . . . . . . . . . . 48

4.1 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.2 Font sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.3 Math fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.4 TEX Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.5 Index Key Syntax Examples . . . . . . . . . . . . . . . . . . . 57 4.6 Key Names for graphicx Package . . . . . . . . . . . . . . . . 58

Chapter 1

Things You Need to Know

In the first part of this chapter you will get a short overview about the philosophy and history of L^{A}TEX 2ε. The second part of the chapter focuses on the basic structures of a L^{A}TEX document. After reading this chapter you should have a rough knowledge of how L^{A}TEX works. When reading on, this will help you to integrate all the new information into the big picture.

1.1 The Name of the Game

1.1.1 TEX

TEX is a computer program by Donald E. Knuth [2]. It is aimed at typesetting text and mathematical formulae.

TEX is pronounced “Tech,” with a “ch” as in the German word “Ach” or in Scottish “Loch.” In an ASCII environment TEX becomes TeX.

1.1.2 L^{A}TEX

L^{A}TEX is a macro package which enables authors to typeset and print their work at the highest typographical quality, using a predefined, professional layout. L^{A}TEX was originally written by Leslie Lamport [1]. It uses the TEX formatter as its typesetting engine.

Recently the L^{A}TEX package has been updated by the L^{A}TEX3 team, led by Frank Mittelbach, to include some long-requested improvements and to reunify all the patched versions which have cropped up since the release of L^{A}TEX 2.09 some years ago. To distinguish the new version from the old, it is called L^{A}TEX 2ε. This documentation deals with L^{A}TEX 2ε.

L^{A}TEX is pronounced “Lay-tech” or “Lah-tech.” If you refer to L^{A}TEX in an ASCII environment you type LaTeX. L^{A}TEX 2ε is pronounced “Lay-tech two e” and typed LaTeX2e.

1.2 Basics

1.2.1 Author, Book Designer, and Typesetter

To publish something, authors give their typed manuscript to a publishing company. A book designer of the publishing company then decides the layout of the document (column width, fonts, space before and after headings, …). The book designer writes his instructions into the manuscript and then gives it to a typesetter, who typesets the book according to these instructions.

A human book designer tries to find out what the author had in mind while writing the manuscript. He decides on chapter headings, citations, examples, formulae, etc. based on his professional knowledge and from the contents of the manuscript.

In a L^{A}TEX environment, L^{A}TEX takes the role of the book designer and uses TEX as its typesetter. But L^{A}TEX is “only” a program and therefore needs more guidance. The author has to provide additional information which describes the logical structure of his work. This information is written into the text as “L^{A}TEX commands.”

This is quite different from the WYSIWYG^{1} approach which most modern word processors such as Word for Windows or WordPerfect take. With these applications, authors specify the document layout interactively while typing text into the computer. All along the way, they can see on the screen how the final work will look when it is printed.

When using L^{A}TEX it is normally not possible to see the final output while typing the text. But the final output can be previewed on the screen after processing the file with L^{A}TEX. Then corrections can be made before actually sending the document to the printer.

1.2.2 Layout Design

Typographical design is a craft. Unskilled authors often commit serious formatting errors by assuming that book design is mostly a question of aesthetics—“If a document looks good artistically it is well designed.” But as a document has to be read and not hung up in a picture gallery, the readability and understandability is of much greater importance than the beautiful look of it. Examples:

• The font size and numbering of headings have to be chosen to make the structure of chapters and sections clear to the reader.

• The line length has to be short enough that it does not strain the eyes of the reader, while long enough to fill the page beautifully.

1What you see is what you get.

\underline{1.2 Basics 3}

With WYSIWYG systems, authors often generate aesthetically pleasing documents with very little or inconsistent structure. L^{A}TEX prevents such formatting errors by forcing the author to declare the logical structure of his document. L^{A}TEX then chooses the most suitable layout.

1.2.3 Advantages and Disadvantages

A topic often discussed when people from the WYSIWYG world meet L^{A}TEX people, is “the advantages of L^{A}TEX over a normal word processor” or the opposite. The best thing you can do when such a discussion starts, is to keep a low profile, as it often gets out of hand. But sometimes you cannot escape …

The main advantages of L^{A}TEX over normal word processors are the following:

• Professionally crafted layouts are available which make a document really look as if “printed.”

• The typesetting of mathematical formulae is supported in a convenient way.

• The user only needs to learn a few easy to understand commands, which specify the logical structure of a document. They almost never need to tinker with the actual layout of the document.

• Even complex structures such as footnotes, references, table of contents, and bibliographies can be generated easily.

• For many typographical tasks not directly supported by basic L^{A}TEX, there exist free add-on packages. For example, packages are available to include PostScript graphics or to typeset bibliographies conforming to exact standards. Many of these add-on packages are described in The L^{A}TEX Companion [3].

• L^{A}TEX encourages authors to write well structured texts because this is how L^{A}TEX works—by specifying structure.

• TEX, the formatting engine of L^{A}TEX 2ε, is highly portable and free. Therefore the system runs on almost any hardware platform available.

L^{A}TEX also has some disadvantages:

• More resources (memory, disk-space, computing power) are required to run a L^{A}TEX system than a simple word processor. But things are getting better, as Word for Windows 6.0 needs even more disk space than a normal L^{A}TEX system. When it comes down to processor usage, L^{A}TEX beats any WYSIWYG system, as it only needs a lot of CPU time when a document is actually processed, while WYSIWYG packages hog the CPU all the time.

• Although within a predefined document layout some parameters can be adjusted, the design of a whole new layout is difficult and takes a lot of time.^{2}

1.3 L^{A}TEX Input Files

The input for L^{A}TEX is a plain ASCII text file. You can create it with any text editor. It contains the text of the document as well as the commands which tell L^{A}TEX how to typeset the text.

1.3.1 Spaces

“Whitespace” characters such as blank or tab are treated uniformly as “space” by L^{A}TEX. Several consecutive whitespace characters are treated as one “space”. Whitespace at the start of a line is generally ignored and a single linebreak is treated like a “space”.

An empty line between two lines of text defines the end of a paragraph. Several empty lines are treated the same as one empty line. The text below is an example. On the right hand side is the text from the input file and on the left hand side is the formatted output.

It does not matter whether you enter one or several spaces after a word. An empty line starts a new paragraph.

                                  It does not matter whether you
                                  enter one or several     spaces
                                  after a word.
                                  An empty line starts a new
                                  paragraph.

1.3.2 Special Characters

The following symbols are reserved characters, that either have a special meaning under L^{A}TEX or are not available in all the fonts. If you enter them in your text directly, they will normally not print, but rather coerce L^{A}TEX to do things you did not intend.

     $ & % # _{ }  ~  ^\

As you will see, these characters can be used in your documents all the same by adding a prefix backslash:

2Rumour says, that this is one of the key elements which will be addressed in the upcoming L^{A}TEX3 system

$ & % # { }  \$ \& \% \# \_\{ \}

The other symbols and many more can be printed with special commands in mathematical formulae or as accents.

1.3.3 L^{A}TEX Commands

L^{A}TEX commands are case sensitive and take one of the following two formats:

• They start with a backslash \ and then have a name consisting only of letters. Command names are terminated by a space, a number or any other ‘non-letter’.

• They consist of a backslash and exactly one special character.

L^{A}TEX ignores whitespace after commands. If you want to get a space after a command, you have to put either {} and a blank or a special spacing command after the command name. The {} stops L^{A}TEX from eating up all the space after the command name.

I read that Knuth divides the people working with TEX into TEXnicians and TEXperts. Today is April 1, 1996.

                                  I read that Knuth divides the
                                  people working with \TeX{} into
                                  \TeX{}nicians and \TeX perts.\\
                                  Today is \today.

Some commands need a parameter which has to be given between curly braces { } after the command name. Some commands support optional parameters which are added after the command name in square brackets [ ]. The next example uses some L^{A}TEX commands. Don’t worry about them, they will be explained later.

You can lean on me!  You can \textsl{lean} on me!

Please, start a new line right here! Thank you!

                                  Please, start a new line
                                  right here!\linebreak[3]
                                  Thank you!

1.3.4 Comments

When L^{A}TEX encounters a % character while processing an input file, it ignores the rest of the present line. This is useful for adding notes to the input file, which will not show up in the printed version.

This is an example.

                                  This is an % stupid
                                  % Better: instructive <----
                                  example.

1.4 Input File Structure

When L^{A}TEX 2ε processes an input file it expects it to follow a certain structure. Thus every input file must start with the command

     \documentclass{...}

This specifies what sort of document you intend to write. After that, you can include commands which influence the style of the whole document or you can load packages which add new features to the L^{A}TEX system. To load such a package you use the command

     \usepackage{...}

When all the setup work is done^{3}, you start the body of the text with the command

     \begin{document}

Now you enter the text mixed with some useful L^{A}TEX commands. At the end of the document you add the

     \end{document}

command, which tells L^{A}TEX to call it a day. Anything which follows this command will be ignored by L^{A}TEX.

Figure 1.1 shows the contents of a minimal L^{A}TEX 2ε file. A slightly more complicated input file is given in Figure 1.2.

3The area between \documentclass and \begin{document} is called preamble.

                \documentclass{article}
                \begin{document}
                Small is beautiful.
                \end{document}

Figure 1.1: A Minimal L^{A}TEX File

1.5 The Layout of the Document

1.5.1 Document Classes

The first information L^{A}TEX needs to know when processing an input file is the type of document the author wants to create. This is specified with the

\documentclass command.
 \documentclass[options]{class}

Here class specifies the type of document to be created. Table 1.1 lists the document classes explained in this introduction. The L^{A}TEX 2ε distribution provides additional classes for other documents including letters and slides. The options parameter customises the behaviour of the document class. The options have to be separated by commas. In Table 1.2 the most common options for the standard document classes are listed.

Example: An input file for a L^{A}TEX document could start with the line

     \documentclass[11pt,twoside,a4paper]{article}

it instructs L^{A}TEX to typeset the document as an article with a base font size of eleven points and to produce a layout suitable for double sided printing on a4 paper.

1.5.2 Packages

While writing your document, you will probably find that there are some areas where basic L^{A}TEX cannot solve your problem. If you want to include

       \documentclass[a4paper,11pt]{article}
       \usepackage{latexsym}
       \author{H.~Partl}
       \title{Minimalism}
       \frenchspacing
       \begin{document}
       \maketitle
       \tableofcontents
       \section{Start}
       Well and here begins my lovely article.
       \section{End}
       \ldots{} and here it ends.
       \end{document}

Figure 1.2: Example of a Realistic Journal Article

Table 1.1: Document Classes

article for articles in scientific journals, presentations, short reports, program documentation, invitations, …

report for longer reports containing several chapters, small books, PhD theses, …

book for real books

slides for slides. The class uses big sans serif letters.

Table 1.2: Document Class Options

10pt, 11pt, 12pt Sets the size of the main font for the document. If no option is specified, 10pt is assumed.

a4paper, letterpaper, … Defines the paper size. The default size

is letterpaper. Besides that, a5paper, b5paper,
executivepaper, and legalpaper can be specified.

fleqn Typesets displayed formulae left-aligned instead of centred.

leqno Places the numbering of formulae on the left hand side instead of the right.

titlepage, notitlepage Specifies whether a new page should be started after the document title or not. The article class does not start a new page by default, while report and book do.

twocolumn Instructs L^{A}TEX to typeset the document in two columns.

twoside, oneside Specifies whether double or single sided output should be generated. The classes article and report are single sided and the book class is double sided by default.

openright, openany Makes chapters begin either only on right hand pages or on the next page available. This does not work with the article class, as it does not know about chapters. The report class by default starts chapters on the next page available and the book class starts them on right hand pages. graphics, coloured text or source code from a file into your document, you need to enhance the capabilities of L^{A}TEX. Such enhancements are called packages. Packages are activated with the

 \usepackage[options]{package}

command. Where package is the name of the package and options is a list of keywords which trigger special features in the package. Some packages come with the L^{A}TEX 2ε base distribution (See Table 1.3). Others are provided separately. You may find more information on the packages installed at your site in your Local Guide [4]. The prime source for information about L^{A}TEX is The L^{A}TEX Companion [3]. It contains descriptions of hundreds of packages along with information of how to write your own extensions to L^{A}TEX 2ε.

Table 1.3: Some of the Packages Distributed with L^{A}TEX

doc Allows the documentation of L^{A}TEX programs.

Described in doc.dtx^{a} and in The L^{A}TEX Companion [3].

exscale Provides scaled versions of the maths extension font.

Described in ltexscale.dtx.

fontenc Specifies which font encoding L^{A}TEX should use.

Described in ltoutenc.dtx.

ifthen Provides commands of the form ‘if… then do… otherwise do. …’ Described in ifthen.dtx and The L^{A}TEX Companion [3].

latexsym To access the L^{A}TEX symbol font, you should use the latexsym package. Described in latexsym.dtx and in The L^{A}TEX Companion [3]

makeidx Provides commands for producing indexes. Described in section 4.5 and in The L^{A}TEX Companion [3].

syntonly Processes a document without typesetting it. Described in syntonly.dtx and in The L^{A}TEX Companion [3]. This is useful for quick error checking.

inputenc Allows the specification of an input encoding such as ASCII, ISO Latin-1, ISO Latin-2, 437/850 IBM code pages, Apple Macintosh, Next, ANSI-Windows or user-defined one.

Described in inputenc.dtx.

aThis file should be installed on your system, and you should be able to get a dvi file by typing latex doc.dtx. The same is true for all the other files mentioned in this table.

1.5.3 Page Styles

L^{A}TEX supports three predefined header/footer combinations—so-called page styles. The style parameter of the

 \pagestyle{style}

command defines which one to use. Table 1.4 lists the predefined page styles.

Table 1.4: The Predefined Page Styles of L^{A}TEX

plain prints the page numbers on the bottom of the page in the middle of the footer. This is the default page style.

headings prints the current chapter heading and the page number in the header on each page while the footer remains empty. (This is the style used in this document)

empty sets both the header and the footer to be empty.

It is possible to change the page style of the current page with the command

 \thispagestyle{style}

In The L^{A}TEX Companion [3] there is a description how to create your own headers and footers.

1.6 Big Projects

When working on big documents, you might want to split the input file into several parts. L^{A}TEX has two commands which help you to do that.

 \include{filename}

you can use this command in the document body, to insert the contents of another file. Note that L^{A}TEX will start a new page before processing the material input from filename.

The second command can be used in the preamble. It allows you to instruct L^{A}TEX to only input some of the \included files.

\includeonly{filename,filename,… }

After this command is executed in the preamble of the document. Only \include commands for the filenames which are listed in the argument of the \includeonly command will be executed. Note that there must be no space between the filename and the comma.

The \include command starts typesetting the included text on a new page. This is helpful when you use \includeonly, because the pagebreaks will not move, even when some included files at ommited. Sometimes this might not be desirable. In this case, you can use the

 \input{filename}

command. It simply includes the file specified. No flashy suits, no strings attached.

Typesetting Text

After reading the previous chapter you should know about the basic stuff from which a L^{A}TEX 2ε document is made. In this chapter I will fill in the remaining structure you will need to know in order to produce real world material.

2.1 Linebreaking and Pagebreaking

2.1.1 Justified Paragraphs

Often books are typeset with each line having the same length. L^{A}TEX inserts the necessary linebreaks and spaces between words by optimising the contents of a whole paragraph. If necessary it also hyphenates words that would not fit comfortably on a line. How the paragraphs are typeset depends on the document class. Normally the first line of a paragraph is indented and there is no additional space between two paragraphs. Refer to section 4.2.2 for more information.

In special cases it might be necessary to order L^{A}TEX to break a line:

 \\ or \newline

start a new line without starting a new paragraph.

 \\*

additionally prohibits a pagebreak after the forced linebreak.

 \newpage

starts a new page.

 \linebreak[n], \nolinebreak[n], \pagebreak[n] and \nopagebreak[n]

do what their names say. They enable the author to influence their actions, with the optional argument n. It can be set to a number between zero to four. By setting n to a value below 4 you leave L^{A}TEX the option of ignoring your command if the result would look very bad.

L^{A}TEX always tries to produce the best linebreaks possible. If it cannot find a way to break the lines in a manner which meets its high standards, it lets one line stick out on the right of the paragraph. L^{A}TEX then complains (“overfull hbox”) while processing the input file. This happens most often when L^{A}TEX cannot find a suitable place to hyphenate a word. By giving the \sloppy command you can instruct L^{A}TEX to lower its standards a little. It then prevents such over-long lines by increasing the inter-word spacing — even if the final output is not optimal. In this case a warning (“underfull hbox”) is given to the user. In most cases the result does not look too bad. The command \fussy acts in the opposite direction. Just in case you want to see L^{A}TEX complaining all over the place.

2.1.2 Hyphenation

L^{A}TEX hyphenates words whenever necessary. If the hyphenation algorithm does not find the correct hyphenation points you can remedy the situation by using the following commands, to tell TEX about the exception.

The command

 \hyphenation{word list}

causes the words listed in the argument to be hyphenated only at the points marked by “-”. This command should be given in the preamble of the input file and should only contain words built from normal letters. The case of the letters is ignored. The example below will allow “hyphenation” to be hyphenated as well as “Hyphenation” and it prevents “FORTRAN”, “Fortran” and “fortran” from being hyphenated at all. No special characters or symbols are allowed in the argument.

Example:

     \hyphenation{FORTRAN Hy-phen-a-tion}

The command \- inserts a discretionary hyphen into a word. This also becomes the only point hyphenation is allowed in this word. This command is especially useful for words containing special characters (e.g. accented characters), because L^{A}TEX does not automatically hyphenate words containing special characters.

I think this is: supercalifragilisticexpialidocious

                                  I think this is: su\-per\-cal\-%
                                  i\-frag\-i\-lis\-tic\-ex\-pi\-%
                                  al\-i\-do\-cious

Several words can be kept together on one line with the command

 \mbox{text}

It causes its argument be kept together under all circumstances.

My phone number will change soon. It will be 0116 291 2319.

The parameter filename should contain the name of the file.

                                  My phone number will change soon.
                                  It will be \mbox{0116 291 2319}.
                                  The parameter
                                  \mbox{\emph{filename}} should
                                  contain the name of the file.

2.2 Special Characters and Symbols

2.2.1 Quotation Marks

For quotation marks you should not use the " as on a typewriter. In publishing there are special opening and closing quotation marks. In L^{A}TEX, use two `s on for opening quotation marks and two 's for closing quotation marks.

“Please press the ‘x’ key.”  ``Please press the `x' key.''

2.2.2 Dashes and Hyphens

L^{A}TEX knows four kinds of dashes. You can access three of these with different numbers of consecutive dashes. The fourth sign is actually no dash at all: It is the mathematical minus sign:

daughter-in-law, X-rated pages 13–67 yes—or no? 0, 1 and −1

                                  daughter-in-law, X-rated\\
                                  pages 13--67\\
                                  yes---or no? \\
                                  $0$, $1$ and $-1$

The names for these dashes are: - hyphen, -- en-dash, --- em-dash and $-$ minus sign.

2.2.3 Ellipsis ( …)

On a typewriter a comma or a period takes the same amount of space as any other letter. In book printing these characters occupy only a little space and are set very close to the preceding letter. Therefore you cannot enter “ellipsis” by just typing three dots, as the spacing would be wrong. Instead there is a special command for these dots. It is called

 \ldots

Not like that ... but like that: New York, Tokyo, Budapest, …

                                  Not like that ... but like that:\\
                                  New York, Tokyo, Budapest, \ldots

2.2.4 Ligatures

Some letter combinations are typeset not just by setting the different letters one after the other, but actually by using special symbols.

ff fi fl ffi… instead of ff fi fl ffi …

These so-called ligatures can be prohibited by inserting a \mbox{} between the two letters in question. This might be necessary with words built from two words.

Not shelfful but shelfful

                                  Not shelfful\\
                                  but shelf\mbox{}ful

2.2.5 Accents and Special Characters

L^{A}TEX supports the use of accents and special characters from many languages. Table 2.1 shows all sorts of accents being applied to the letter o. Naturally other letters work too.

To place an accent on top of an i or a j, their dots have to be removed. This is accomplished by typing \i and \j.

Hôtel, naïve, élève, smørrebrød, ¡Señorita!, Schönbrunner Schloß Straße

                                  H\^otel, na\"\i ve, \'el\`eve,\\
                                  sm\o rrebr\o d, !`Se\~norita!,\\
                                  Sch\"onbrunner Schlo\ss{}
                                  Stra\ss e

2.3 International Language Support

If you need to write documents in languages other than English, L^{A}TEX must apply different hyphenation rules in order to produce correct output.

For many languages, these changes can be accomplished by using the babel package by Johannes Braams. To use this package, your L^{A}TEX system has to be specially configured. Your Local Guide [4] should give more information on this.

If your system is already appropriately configured, you can activate the babel package by adding the command

 \usepackage[language]{babel}

after the \documentclass command. Which languages your system supports should also be listed in the Local Guide.

For some languages babel also specifies new commands, which simplify the input of special characters. The German language for example, contains a lot of umlauts (äöü). With babel you can enter an ö by typing "o instead

of \"o.

Some computer systems allow you to input special characters directly from the keyboard. L^{A}TEX can handle such characters. Since the December 1994 release of L^{A}TEX 2ε, support for several input encodings is included in the basic distribution of L^{A}TEX 2ε. Check the inputenc package. When using this package you should consider that other people might not be able to display your input files on their computer, because they use a different encoding. For example, the German umlaut ä on a PC is encoded as 132 and on some Unix systems using ISO-LATIN 1 it is encoded as 228. Therefore, use this feature with care.

Table 2.1: Accents and Special Characters

        ò \`o  ó \'o  ô \^o  õ \~o
        ō \=o  ȯ \.o  ö \"o
        ŏ \u o ǒ \v o ő \H o  o̧ \c o
        o. \d o   ō \b o oo \t oo
        œ \oe  Œ \OE  æ \ae  Æ \AE
        å \aa  Å \AA
        ø \o  Ø \O  ł \l  Ł \L
        ı \i  ȷ \j  ¡ !`  ¿ ?`

2.4 The Space between Words

To get a straight right margin in the output, L^{A}TEX inserts varying amounts of space between the words. At the end of a sentence it inserts slightly more space, as this makes the text more readable. L^{A}TEX assumes that sentences end with periods, question marks or exclamation marks. If a period follows an uppercase letter this is not taken as a sentence ending since periods after uppercase letters are normally for abbreviations.

Any exception from these assumptions has to be specified by the author. A backslash in front of a space generates a space which will not be enlarged. A tilde ‘~’ character generates a space which cannot be enlarged and which additionally prohibits a linebreak. The command \@ in front of a period specifies, that this period terminates a sentence even when it follows a uppercase letter.

Mr. Smith was happy to see her cf. Fig. 5 I like BASIC. What about you?

                                  Mr.~Smith was happy to see her\\
                                  cf.~Fig.~5\\
                                  I like BASIC\@. What about you?

The additional space after periods can be disabled with the command

 \frenchspacing

which tells L^{A}TEX not to insert any more space after a period than after ordinary character. This is very common in non-English languages, except for bibliographies. If you use \frenchspacing, the command \@ is not necessary.

2.5 Titles, Chapters, and Sections

To help the reader find his or her way through your work, you should divide it into chapters, sections, and subsections. L^{A}TEX supports this with special commands which take the section title as their argument. It is up to you to use them in the correct order.

For the article class the following sectioning commands are available:

     \section{...}           \paragraph{...}
     \subsection{...}        \subparagraph{...}
     \subsubsection{...}     \appendix

For the report and the book class you can use two additional sectioning commands:

     \part{...}              \chapter{...}

As the article class does not know about chapters, it is quite easy to add articles as chapters to a book. The spacing between sections, the numbering and the font size of the titles will be set automatically by L^{A}TEX.

Two of the sectioning commands are a bit special:

• The \part command does not influence the numbering sequence of chapters.

• The \appendix command does not take an argument. It just changes the chapter^{1} numbering to letters.

L^{A}TEX creates a table of contents by taking the section headings and page numbers from the previous run of the document. The command

 \tableofcontents

expands to a table of contents at the place where it is issued. A new document has to be processed (“L^{A}TEXed”) twice to get a correct table of contents. In some circumstances it might be necessary to compile the document a third time. L^{A}TEX will tell you when this is necessary.

All sectioning commands listed above also exist as “starred” versions. A “starred” version of a command is built by adding a star * after the command name. They generate section headings which will not show up in the table of contents and which will not get numbered. The command

\section{Help} for example would become \section*{Help}.

Normally the section headings show up in the table of contents exactly as they were entered in the text. Sometimes this is not possible, because the heading is too long to fit into the table of contents. The entry for the table of contents can therefore be specified as an optional argument before the actual heading.

     \chapter[Read it! It's Exciting]{This is a very long
                      and especially boring title}

The title of the whole document is generated by issuing a

 \maketitle

command. The contents of the title has to be defined by the commands

 \title{...}, \author{...} and optionally \date{...}

before calling \maketitle. In the argument of \authors you can supply several names separated by \and commands.

1For the article style it changes the section numbering

An example of some of the above mentioned commands can be found in Figure 1.2 on page 7.

Apart from the sectioning commands explained above, L^{A}TEX 2ε introduced 3 additional commands, for use with the book class.

 \frontmatter, \mainmatter and \backmatter

They are useful for dividing your publication. The commands alter chapter headings, and page numbering to work as you would expect it in a book.

2.6 Cross References

In books, reports and articles there are often cross references to figures, tables and special segments of text. L^{A}TEX provides the following commands for cross referencing

 \label{marker}, \ref{marker} and \pageref{marker}

Where marker is an identifier chosen by the user. L^{A}TEX replaces \ref by the number of the section, subsection, figure, table, or theorem where the corresponding \label command was issued. \pageref prints the page number of the corresponding \label command. Here also the numbers from the previous run are used.

A reference to this subsection looks like: “see section 2.6 on page 20.”

                                  A reference to this subsection
                                  \label{sec:this} looks like:
                                  ``see section~\ref{sec:this} on
                                  page~\pageref{sec:this}.''

2.7 Footnotes

With the command

a footnote will be printed at the foot of the current page.

                                  Footnotes\footnote{This
                                    is a footnote} are often used
                                  by people using \LaTeX.

\footnote{footnote text} Footnotes^{a} are often used by people using L^{A}TEX.

aThis is a footnote

2.8 Emphasized Words

In manuscripts produced by typewriter, important words get \underline{underlined}. In printed books these words are emphasised. The command to switch to an emphasised font is called

 \emph{text}

Its argument is the text to be emphasised.

If you use emphasising in an already emphasised text, then L^{A}TEX uses an upright font for emphasising.

                                  \emph{If you use
                                    \emph{emphasising} in an
                                    already emphasised text, then
                                    \LaTeX{} uses an
                                    \emph{upright} font for
                                    emphasising.}

2.9 Environments

To typeset special purpose text, L^{A}TEX defines many different environments for all sorts of formatting:

 \begin{name} text \end{name}

Where name is the name of the environment. Environments can be called several times within each other as long as the calling order is maintained.

     \begin{aaa}...\begin{bbb}...\end{bbb}...\end{aaa}

In the following sections all important environments are explained.

2.9.1 Itemise, Enumerate, and Description

The itemize environment is suitable for simple lists, the enumerate environment for enumerated lists, and the description environment for descriptions.

22  Typesetting Text

1. You can mix the list environments to your taste:

• But it might start to look silly.

- With a dash.

2. Therefore remember:

Stupid things will not become smart because they are in a list.

Smart things though, can be presented beautifully in a list.

\begin{enumerate}
\item You can mix the list
environments to your taste:
\begin{itemize}
\item But it might start to
look silly.
\item[-] With a dash.
\end{itemize}
\item Therefore remember:
\begin{description}
\item[Stupid] things will not
become smart because they are
in a list.
\item[Smart] things though, can be
presented beautifully in a list.
\end{description}
\end{enumerate}

2.9.2 Flushleft, Flushright, and Center

The environments flushleft and flushright generate paragraphs which are either left or right aligned. The center environment generates centred text. If you do not issue \\ to specify linebreaks, L^{A}TEX will automatically determine linebreaks.

This text is left aligned. L^{A}TEX is not trying to make each line the same length.

\begin{flushleft}
This text is\\ left aligned.
\LaTeX{} is not trying to make
each line the same length.
\end{flushleft}

This text is right

aligned. L^{A}TEX is not trying to make each line the same length.

\begin{flushright}
This text is right\\ aligned.
\LaTeX{} is not trying to make
each line the same length.
\end{flushright}

At the centre of the earth

\begin{center}
At the centre\\of the earth
\end{center}

2.9.3 Quote, Quotation, and Verse

The quote environment is useful for quotes, important phrases and examples.

2.9 Environments  23

A typographical rule of thumb for the line length is:

No line should contain more than 66 characters. This is why L^{A}TEXpages have such large borders by default.

That’s why multicolumn print is often used in newspapers.

A typographical rule of thumb
for the line length is:
\begin{quote}
No line should contain more than
66~characters.
This is why \LaTeX pages have
such large borders by default.
\end{quote}
That's why multicolumn print is
often used in newspapers.

There are two similar environments: the quotation and the verse environments. The quotation environment is useful for longer quotes going over several paragraphs, because it does indent paragraphs. The verse environment is useful for poems where the line breaks are important. The lines are separated by issuing a \\ at the end of a line and a empty line after each verse.

I know only one English poem by heart. It is about Humpty Dumpty.

Humpty Dumpty sat on

a wall: Humpty Dumpty had a

great fall. All the King’s horses

and all the King’s men

Couldn’t put Humpty together again.

I know only one English poem by
heart. It is about Humpty Dumpty.
\begin{flushleft}
\begin{verse}
Humpty Dumpty sat on a wall:\\
Humpty Dumpty had a great fall.\\
All the King's horses and all
the King's men\\
Couldn't put Humpty together
again.
\end{verse}
\end{flushleft}

2.9.4 Printing Verbatim

Text which is enclosed between \begin{verbatim} and \end{verbatim}

will be directly printed, as if it was typed on a typewriter, with all linebreaks and spaces, without any L^{A}TEX command being executed.

Within a paragraph, similar functionality can be accessed with

\verb+text+

The + is just an example delimiter character. You can use any character except letters, * or blank. Many L^{A}TEX examples in this booklet are typeset with this command.

The \ldots command …

10 PRINT "HELLO WORLD ";
20 GOTO 10

                                  The \verb|\ldots| command \ldots
                                  \begin{verbatim}
                                  10 PRINT "HELLO WORLD ";
                                  20 GOTO 10
                                  \end{verbatim}

the␣starred␣version␣of
the␣␣␣␣␣␣verbatim
environment␣emphasises
the␣spaces␣␣␣in␣the␣text

                                  \begin{verbatim*}
                                  the starred version of
                                  the      verbatim
                                  environment emphasises
                                  the spaces   in the text
                                  \end{verbatim*}

The \verb command can be used in a similar fashion with a star:

like␣␣␣this␣:-)␣                  \verb*|like   this :-) |

The verbatim environment and the \verb command may not be used within parameters of other commands.

2.9.5 Tabular

The tabular environment can be used to typeset beautiful tables with optional horizontal and vertical lines. L^{A}TEX determines the width of the columns automatically.

The table spec argument of the

 \begin{tabular}{table spec}

command defines the format of the table. Use an l for a column of left aligned text, r for right aligned text and c for centred text, p{width} for a column containing justified text with linebreaks, and | for a vertical line.

Within a tabular environment & jumps to the next column, \\ starts a new line and \hline inserts an horizontal line.

                                  \begin{tabular}{|r|l|}
                                  \hline
                                  7C0 & hexadecimal \\
                                  3700 & octal \\
                                  11111000000 & binary \\
                                  \hline \hline
                                  1984 & decimal \\
                                  \hline
                                  \end{tabular}

7C0 hexadecimal 3700 octal 11111000000 binary

1984 decimal

Welcome to Boxy’s paragraph. We sincerely hope you’ll all enjoy the show.

                                  \begin{tabular}{|p{4.7cm}|}
                                  \hline
                                  Welcome to Boxy's paragraph.
                                  We sincerely hope you'll
                                  all enjoy the show.\\
                                  \hline
                                  \end{tabular}

With the @{...} construct it is possible to specify the column separator. This command kills the intercolumn space and replaces it with whatever is included in the curly braces. One common use for this command is explained below in the decimal alignment problem. Another possible usage is to suppress leading space in a table with @{}.

no leading space

                                  \begin{tabular}{@{} l @{}}
                                  \hline
                                  no leading space\\\hline
                                  \end{tabular}

leading space left and right

                                  \begin{tabular}{l}
                                  \hline
                                  leading space left and right\\
                                  \hline
                                  \end{tabular}

Since there is no built-in way to align numeric columns on a decimal point^{2}, we can “cheat” and do it by using two columns: a right-aligned integer and a left-aligned fraction. The @{} command in the \begin{tabular} line replaces the normal intercolumn spacing with just a “.”, giving the appearance of a single, decimal-point-justified column. Don’t forget to replace the decimal point in your numbers with a column separator (&)! A column label can be placed above our numeric “column” by using the \multicolumn command.

Pi expression  Value

                                  \begin{tabular}{c r @{.} l}
                                  Pi expression       &
                                  \multicolumn{2}{c}{Value} \\
                                  \hline
                                  $\pi$               & 3&1416  \\
                                  $\pi^{\pi}$         & 36&46   \\
                                  $(\pi^{\pi})^{\pi}$ & 80662&7 \\
                                  \end{tabular}

π  3.1416 π^{π}  36.46 (π^{π})^{π}  80662.7 2If the ‘tools’ bundle is installed on your system, have a look at the dcolumn package

2.10 Floating Bodies

Today most publications contain a lot of figures and tables. These elements need special treatment because they cannot be broken across pages. One method would be to start a new page every time a figure or a table is too large to fit on the present page. This approach would leave pages partially empty which looks very bad.

The solution to this problem is to ‘float’ any figure or table, which does not fit on the current page, to a later page while filling the current page with body text. L^{A}TEX offers two environments for floating bodies. One for tables and one for figures. To take full advantage of these two environments it is important to understand approximately how L^{A}TEX handles floats internally. Otherwise floats may become a major source of frustration because L^{A}TEX never puts them where you want them to be.

Let’s first have a look at the commands L^{A}TEX supplies for floats:

Any material enclosed in a figure or table environment will be treated as floating matter. Both float environments support an optional parameter

\begin{figure}[placement specifier] or \begin{table}[placement specifier]

called the placement specifier. This parameter is used to tell L^{A}TEX about the locations the float is allowed to be moved to. A placement specifier is constructed by building a string of float placing permissions. See Table 2.2.

A table could be started with the following line e.g.

     \begin{table}[!hbp]

The placement specifier [!hbp] allows L^{A}TEX to place the table right here (h) or at the bottom (b) of some page or on a special floats page (p) and all that even if it does not look that good (!). If no placement specifier is given, the standard classes assume [tbp].

Table 2.2: Float Placing Permissions

Spec Permission to place the float …

h here at the very place in the text where it occurred. This is

useful mainly for small floats.

t at the top of a page
b at the bottom of a page
p on a special page containing only floats.

! without considering most of the internal parameters^{a} which could stop this float from being placed.

aSuch as the maximum number of floats allowed on one page

L^{A}TEX will place every float it encounters, according to the placement specifier supplied by the author. If a float cannot be placed on the current page it is deferred either to the figures or the tables queue^{3}. When a new page is started, L^{A}TEX first checks if it is possible to fill a special ‘float’ page with floats from the queues. If this is not possible, the first float on each queue is treated as if they had just occurred in the text: L^{A}TEX tries again to place them according to their respective placement specifiers (except ‘h’ which is no longer possible). Any new floats occurring in the text get placed into the appropriate queues. L^{A}TEX strictly maintains the original order of appearance for each type of float. That’s why a figure which cannot be placed, pushes all the further figures to the end of the document. Therefore:

If L^{A}TEX is not placing the floats as you expected, it is often only one float jamming one of the two float queues.

Having explained the difficult bit, there are some more things to mention about the table and figure environments. With the

\caption{caption text}

command you can define a caption for the float. A running number and the string “Figure” or “Table” will be added by L^{A}TEX.

The two commands

 \listoffigures and \listoftables

operate analogously to the \tableofcontents command, printing a list of figures or tables respectively. In these lists, the whole caption will be repeated. If you tend to use long captions, you must have a shorter version of the caption going into the lists. This is accomplished by entering the short version in brackets after the \caption command.

     \caption[Short]{LLLLLoooooonnnnnggggg}

With \label and \ref you can create a reference to a float within your text.

The following example draws a square and inserts it into the document. You could use this if you wanted to reserve space for images you are going to paste into the finished document.

     Figure~\ref{white} is an example of Pop-Art.
     \begin{figure}[!hbp]
     \makebox[\textwidth]{\framebox[5cm]{\rule{0pt}{5cm}}}
     \caption{Five by Five in Centimetres} \label{white}
     \end{figure}

3These are fifo - ‘first in first out’ queues!

In the example above^{4} L^{A}TEX will try really hard (!) to place the figure right here (h). If this is not possible, it tries to place the figure at the bottom (b) of the page. Failing to place the figure on the current page, it determines if it is possible to create a float page containing this figure and maybe some tables from the tables queue. If there is not enough material for a special float page, L^{A}TEX starts a new page and once more treats the figure as if it had just occurred in the text.

Under certain circumstances it might be necessary to use the

 \clearpage or even the \cleardoublepage

command. It orders L^{A}TEX to immediately place all floats remaining in the queues and then start a new page. \cleardoublepage goes to a new lefthand page on top of that.

Later in this introduction you will learn how to include PostScript drawings into your L^{A}TEX 2ε documents.

2.11 Adding New Commands and Environments

In the first chapter, I explained that L^{A}TEX needs information about the logical structure of the text to pick the right layout. This is a neat idea, but in practice you often get to the limits of this, because L^{A}TEX just does not have a specialised environment or command for your exact purpose.

One solution is to use several L^{A}TEX commands to produce the layout you have in mind. If you have to do this once, then there is no problem, but if it occurs repeatedly, it takes a lot of time. If you ever want to change the layout you have to work through the whole input file and edit all the items in question.

To solve this problem, L^{A}TEX allows you to define your own commands and environments to the language.

2.11.1 New Commands

To add your own commands, use the

 \newcommand{name}[num]{definition}

command. Basically, the command requires two arguments. The name of the command you want to create and the definition of the command. The num argument in square brackets is optional. You can use it to create new commands which themselves take up to 9 arguments.

4assuming the figure queue is empty

The following two examples should help you to get the idea. The first example defines a new command called \tnss this is short for “The Not So Short Introduction to L^{A}TEX 2ε”. Such a command could come in handy if you have to write the title of this book over and over again.

“The not so Short Introduction to L^{A}TEX 2ε” … “The not so Short Introduction to L^{A}TEX 2ε”

                                  \newcommand{\tnss}
                                      {The not so Short Introduction
                                          to \LaTeXe}
                                  % in the document body :
                                  ``\tnss'' \ldots{} ``\tnss''

The next example illustrates how to use the num argument. The #1 sequence gets replace by the argument you specify. If you wanted to use more than one argument, use #2 and so on.

• The not so Short Introduction to L^{A}TEX 2ε

• The very Short Introduction to L^{A}TEX 2ε

                                  \newcommand{\txsit}[1]
                                      {The \emph{#1} Short
                                        Introduction to \LaTeXe}
                                  % in the document body:
                                  \begin{itemize}
                                  \item \txsit{not so}
                                  \item \txsit{very}
                                  \end{itemize}

L^{A}TEX will not allow you to create a new command which already exists. If you explicitly want to override an existing command you have to use the \renewcommand. Apart from its name, it uses the same syntax as the \newcommand command. In certain cases you might also want to use the \providecommand command. It works like \newcommand, but if the new command is already defined, L^{A}TEX 2ε just ignores it.

2.11.2 New Environments

Analogously to the \newcommand command there is a command to create your own environments. When writing this introduction, I have created special environments for the structures I used repeatedly through out the introduction: “examples”, “bits of code” and “command definition boxes”. The \newenvironment command uses the following syntax:

 \newenvironment{name}[num]{before}{after}

Like the \newcommand command, you can use \newenvironment with an optional argument and without. The material specified in the before argument is processed before the text in the environment gets processed. The material in the after argument gets processed when the \end{name} command is encountered.

The example below illustrates the usage of the \newenvironment command.

My humble subjects …

                                  \newenvironment{king}
                                      {\begin{quote}}{\end{quote}}
                                  % use this in the body part
                                  \begin{king}
                                  My humble subjects \ldots
                                  \end{king}

The num argument is used the same way as in the \newcommand command. L^{A}TEX makes sure that you do not define an environment which already exists. If you ever want to change an existing command, you can use the \renewenvironment command. It uses the same syntax as the

\newenvironment command.

Typesetting Mathematical Formulae

Now you are ready! In this chapter we will attack the main strength of TEX: mathematical typesetting. But be warned, this chapter only scratches the surface. While the things explained here are sufficient for many people, don’t despair if you can’t find a solution to your mathematical typesetting needs. It is highly likely that your problem is addressed in AMS-L^{A}TEX ^{1} or some other package.

3.1 General

L^{A}TEX has a special mode for typesetting mathematics. Mathematical text within a paragraph is entered between \( and \), between $ and $ or between

\begin{math} and \end{math}.

Add a squared and b squared to get c squared. Or using a more mathematical approach: c^{2} = a^{2} + b^{2}

                                  Add $a$ squared and $b$ squared
                                  to get $c$ squared. Or using
                                  a more mathematical approach:
                                  $c^{2}=a^{2}+b^{2}$

TEX is pronounced as τϵχ.

100 m^{3} of water

This comes from my ♡

                                  \TeX{} is pronounced as
                                  $\tau\epsilon\chi$.\\[6pt]
                                  100~m$^{3}$ of water\\[6pt]
                                  This comes from my $\heartsuit$

It is preferable to display larger mathematical equations or formulae, that is to typeset them on separate lines. Therefore you enclose them between \[

and \] or between \begin{displaymath} and \end{displaymath}. This

1CTAN:/tex-archive/macros/latex/packages/amslatex

produces formulae which are not numbered. If you want L^{A}TEX to number them, you can use the equation environment.

Add a squared and b squared to get c squared. Or using a more mathematical approach:

c^{2} = a^{2} + b^{2}

And just one more line.

                                  Add $a$ squared and $b$ squared
                                  to get $c$ squared. Or using
                                  a more mathematical approach:
                                  \begin{displaymath}
                                  c^{2}=a^{2}+b^{2}
                                  \end{displaymath}
                                  And just one more line.

With \label and \ref you can reference an equation within the text.

ϵ > 0  (3.1)

From (3.1) we gather …

                                  \begin{equation} \label{eq:eps}
                                  \epsilon > 0
                                  \end{equation}
                                  From (\ref{eq:eps}) we gather
                                  \ldots

Note that expressions will be typeset if a different style if displayed: contrast

lim_{n→∞} ∑_{k=1}^{n} \frac{1}{k^{2}} = \frac{π^{2}}{6} ; lim_{n→∞} ∑_{k=1}^{n} \frac{1}{k^{2}} = \frac{π^{2}}{6} $\lim_{n \to \infty}

                                  \sum_{k=1}^n \frac{1}{k^2}
                                  = \frac{\pi^2}{6}$
                                  \begin{displaymath}
                                  \lim_{n \to \infty}
                                  \sum_{k=1}^n \frac{1}{k^2}
                                  = \frac{\pi^2}{6}
                                  \end{displaymath}

There are differences between math mode and text mode. For example, in math mode:

1. Most spaces and linebreaks do not have any significance, as all spaces are either derived logically from the mathematical expressions or have to be specified using special commands such as \,, \quad or \qquad.

∀x ∈ R :  x^{2} ≥ 0  (3.2)

                                     \begin{equation}
                                     \forall x \in \mathbf{R}:
                                     \qquad x^{2} \geq 0
                                     \end{equation}

2. Empty lines are not allowed. Only one paragraph per formula.

3. Each letter is considered to be the name of a variable and will be typeset as such. If you want to typeset normal text within a formula (normal upright font and normal spacing) then you have to enter the text using the \textrm{...} commands.

x^{2} ≥ 0 for all x ∈ R (3.3)

                                     \begin{equation}
                                     x^{2} \geq 0\qquad
                                     \textrm{for all }x\in\mathbf{R}
                                     \end{equation}

Mathematicians can be very fussy about which symbols are used: it would be conventional here to use ‘blackboard bold’ which is obtained by \mathbb from the package amsfonts or amssymb. The last example becomes

x^{2} ≥ 0 for all x ∈ ℝ

                                  \begin{displaymath}
                                  x^{2} \geq 0\qquad
                                  \textrm{for all }x\in\mathbb{R}
                                  \end{displaymath}

3.2 Grouping in Math Mode

Most math mode commands act only on the next character. So if you want several characters affected by a command you have to group them together using curly braces: {...}.

a^{x} + y ≠ a^{x+y} (3.4)

                                  \begin{equation}
                                  a^x+y \neq a^{x+y}
                                  \end{equation}

3.3 Building Blocks of a Mathematical Formula

In this section the most important commands used in mathematical typesetting will be described. For a list of all symbols available take a look at section 3.10 on page 42.

Lowercase Greek letters are entered as \alpha, \beta, \gamma, …, uppercase letters^{2} are entered as \Gamma, \Delta, …

λ, ξ, π, μ, Φ, Ω  $\lambda,\xi,\pi,\mu,\Phi,\Omega$

Exponents and Subscripts can be specified using the ^and the _ character.

2There is no uppercase Alpha defined in L^{A}TEX 2ε because it looks the same as a normal roman A. Once the new math coding is done, things will change.

{a_{1}  x^{2}  e^{−αt}  a_{ij}^{3} \atop e^{x^{2}} ≠ e^{x}^{2}}

                                  $a_{1}$ \qquad $x^{2}$ \qquad
                                  $e^{-\alpha t}$ \qquad
                                  $a^{3}_{ij}$\\
                                  $e^{x^2} \neq {e^x}^2$

The square root is entered as \sqrt, the n^{th} root is generated with \sqrt[n]. The size of the root sign is determined automatically by L^{A}TEX. If just the sign is needed use \surd.

\begin{matrix}\sqrt{x} & \sqrt{x^{2} + \sqrt{y}} & ^{3} & \sqrt{2} \\ √[x^{2} + y^{2}] \end{matrix}

                                  $\sqrt{x}$ \qquad
                                  $\sqrt{ x^{2}+\sqrt{y} }$
                                  \qquad $\sqrt[3]{2}$\\[3pt]
                                  $\surd[x^2 + y^2]$

The commands \overline and \underline create horizontal lines directly over or under an expression.

\overline{m + n}  $\overline\{m+n\}$

The commands \overbrace and \underbrace create long horizontal braces over or under an expression.

\underbrace{a + b + · · · + z}_{26}  $\underbrace\{ a+b+\cdots+z \}_\{26\}$

To add mathematical accents such as small arrows or tilde signs to variables you can use the commands given in table 3.1. Wide hats and tildes, covering several characters are generated with \widetilde and \widehat. The ' symbol gives a prime.

y = x^{2}  y′ = 2x  y′′ = 2

                                  \begin{displaymath}
                                  y=x^{2}\qquad y'=2x\qquad y''=2
                                  \end{displaymath}

Often vectors are specified by adding small arrow symbols on top of a variable. This is done with the \vec command. To denote the vector from A to B the two commands \overrightarrow and \overleftarrow are useful.

a⃗ \overrightarrow{AB}

                                  \begin{displaymath}
                                  \vec a\quad\overrightarrow{AB}
                                  \end{displaymath}

Names of log-like functions are often typeset in an upright font and not italic as variables. Therefore L^{A}TEX supplies the following commands to typeset the most important function names:

\arccos   \cos    \csc   \exp   \ker     \limsup  \min   \sinh
\arcsin   \cosh   \deg   \gcd   \lg      \ln      \Pr    \sup
\arctan   \cot    \det   \hom   \lim     \log     \sec   \tan
\arg      \coth   \dim   \inf   \liminf  \max     \sin   \tanh

lim_{n→0} \frac{sin x}{x} = 1

                                  \[\lim_{n \rightarrow 0}
                                  \frac{\sin x}{x}=1\]

For the modulo function there are two commands: \bmod for the binary operator “a mod b” and \pmod for expressions such as “x ≡ a (mod b).”

A built-up fraction is typeset with the \frac{...}{...} command. Often the slashed form 1/2 is preferable, because it looks better for small amounts of ‘fraction material.’

{1\frac{1}{2} hours \atop \frac{x^{2}}{k + 1}  x^{\frac{2}{k+1}}  x^{1/2}}

                                  $1\frac{1}{2}$~hours
                                  \begin{displaymath}
                                  \frac{ x^{2} }{ k+1 }\qquad
                                  x^{ \frac{2}{k+1} }\qquad
                                  x^{ 1/2 }
                                  \end{displaymath}

To typeset binomial coefficients or similar structures you can use either

the command {... \choose ...} or {... \atop ...}. The second com-

mand produces the same output as the first one, but without braces.

\binom{n}{k} {x \atop y + 2}

                                  \begin{displaymath}
                                  {n \choose k}\qquad {x \atop y+2}
                                  \end{displaymath}

The integral operator is generated with \int, the sum operator with \sum. The upper and lower limits are specified with ^and _as with subscripts and superscripts.

∑_{i=1}^{n} ∫0^{\frac{π}{2}}

                                  \begin{displaymath}
                                  \sum_{i=1}^{n} \qquad
                                  \int_{0}^{\frac{\pi}{2}} \qquad
                                  \end{displaymath}

For braces and other delimiters there exist all types of symbols in TEX (e.g. [ ⟨ ∥ ↕). Round and square braces can be entered with the corresponding keys, curly braces with \{, all other delimiters are generated with special commands (e.g. \updownarrow). For a list of all delimiters available, check table 3.8 on page 44.

a, b, c ≠ {a, b, c}

                                  \begin{displaymath}
                                  {a,b,c}\neq\{a,b,c\}
                                  \end{displaymath}

If you put the command \left in front of an opening delimiter or \right in front of a closing delimiter, TEX will automatically determine the correct size of the delimiter. Note, the you must close every \left with a corresponding \right. If you don’t want any thing on the right, use the invisible

‘\right.’!

1 + (\frac{1}{1 − x^{2}} )^{3}

                                  \begin{displaymath}
                                  1 + \left( \frac{1}{ 1-x^{2} }
                                      \right) ^3
                                  \end{displaymath}

In some cases it is necessary to specify the correct size of a mathematical delimiter by hand, therefore you can use the commands \big, \Big, \bigg and \Bigg as prefixes to most delimiter commands^{3}.

\begin{matrix}((x + 1)(x − 1))^{2} \\ ( & ( & ( & ( & \} & \} & \} & \} & ∥ & ∥ & ∥ & ∥ \end{matrix}

                                  $\Big( (x+1) (x-1) \Big) ^{2}$\\
                                  $\big(\Big(\bigg(\Bigg($\quad
                                  $\big\}\Big\}\bigg\}\Bigg\}$\quad
                                  $\big\|\Big\|\bigg\|\Bigg\|$

To enter three dots into a formula you can use several commands. \ldots typesets the dots on the baseline, \cdots sets them centred. Beside that there are the commands \vdots for vertical and \ddots for diagonal dots. In section 3.5 you can find another example.

x_{1}, …, x_{n} x_{1} + ⋯ + x_{n}

                                  \begin{displaymath}
                                  x_{1},\ldots,x_{n} \qquad
                                  x_{1}+\cdots+x_{n}
                                  \end{displaymath}

3.4 Math Spacing

If the spaces within formulae chosen by TEX are not satisfactory, they can be adjusted by inserting special spacing commands. The most important are: \, for a tiny space, \␣ for a medium sized space ( ␣ stands for a “space” character), \quad and \qquad for large spaces and \! which shrinks a space.

3These commands do not work as expected if a size changing command has been used, or the 11pt or 12pt option has been specified. Use the exscale or amsmath packages to correct this behaviour

∫∫_{D} g(x, y) dx dy

instead of

∫ ∫_{D} g(x, y)dxdy

                                  \newcommand{\ud}{\mathrm{d}}
                                  \begin{displaymath}
                                  \int\!\!\!\int_{D} g(x,y)
                                    \, \ud x\, \ud y
                                  \end{displaymath}
                                  instead of
                                  \begin{displaymath}
                                  \int\int_{D} g(x,y)\ud x \ud y
                                  \end{displaymath}

Note that ‘d’ in the differential is conventionally set in roman type.

3.5 Vertically Aligned Material

To typeset arrays, use the array environment. It works somewhat similar to the tabular environment. The \\ command is used to break the lines.

X = ( \begin{matrix}x_{11} & x_{12} & … \\ x_{21} & x_{22} & … \\ ⋮ & ⋮ & ... \end{matrix} )

                                  \begin{displaymath}
                                  \mathbf{X} =
                                  \left( \begin{array}{ccc}
                                  x_{11} & x_{12} & \ldots \\
                                  x_{21} & x_{22} & \ldots \\
                                  \vdots & \vdots & \ddots
                                  \end{array} \right)
                                  \end{displaymath}

The array environment can also be used to typeset expressions which have one big delimiter by using a “.” as a invisible \right delimiter:

y = \{ \begin{matrix}a & if d > c \\ b + x & in the morning \\ l & all day long \end{matrix}

                                  \begin{displaymath}
                                  y = \left\{ \begin{array}{ll}
                                   a & \textrm{if $d>c$}\\
                                   b+x & \textrm{in the morning}\\
                                   l & \textrm{all day long}
                                    \end{array} \right.
                                  \end{displaymath}

For formulae running over several lines or for equation systems you can use the environments eqnarray and eqnarray* instead of equation. In eqnarray each line gets an equation number. In the eqnarray* no line numbers are produced.

The eqnarray and the eqnarray* environments work like a 3-column table of the form {rcl}, where the middle column can be used for the equal sign or the not-equal sign. Or any other sign you see fit. The \\ command breaks the lines.

f(x) = cos x (3.5)
f′(x) = − sin x (3.6)
∫_{0}^{x} f(y)dy = sin x (3.7)

                                  \begin{eqnarray}
                                  f(x) & = & \cos x     \\
                                  f'(x) & = & -\sin x   \\
                                  \int_{0}^{x} f(y)dy &
                                   = & \sin x
                                  \end{eqnarray}

Notice that there is too much space each side of the middle column, the equal signs. This can be reduced by setting \setlength\arraycolsep{2pt} as in the next example.

Long equations will not be automatically divided into neat bits. The author has to specify where to break them and how much to indent. The following two methods are the most common ones used to achieve this.

\begin{aligned}sin x & = & x − \frac{x^{3}}{3!} + \frac{x^{5}}{5!} − \\ & & − \frac{x^{7}}{7!} + ⋯ & (3.8) \end{aligned}

                                  {\setlength\arraycolsep{2pt}
                                  \begin{eqnarray}
                                  \sin x & = & x -\frac{x^{3}}{3!}
                                       +\frac{x^{5}}{5!}-{}
                                                      \nonumber\\
                                   & & {}-\frac{x^{7}}{7!}+{}\cdots
                                  \end{eqnarray}}

\begin{matrix}cos x = 1 − \frac{x^{2}}{2!} + \\ + \frac{x^{4}}{4!} − \frac{x^{6}}{6!} + ⋯ & (3.9) \end{matrix}

                                  \begin{eqnarray}
                                  \lefteqn{ \cos x = 1
                                       -\frac{x^{2}}{2!} +{} }
                                                      \nonumber\\
                                   & & {}+\frac{x^{4}}{4!}
                                       -\frac{x^{6}}{6!}+{}\cdots
                                  \end{eqnarray}

The \nonumber command causes L^{A}TEX not to generate a number for this equation.

It can be difficult to get vertically aligned equations to look right with these methods; the package amsmath provides a more powerful set of alternatives.

3.6 Math Font Size

In math mode TEX selects the font size according to context. Superscripts for example get typeset in a smaller font. If you want to add roman text to an equation and use the \textrm command, the font size switching mechanism will not work, as \textrm temporarily escapes to text mode. Use \mathrm instead to keep the size switching mechanism active. But pay attention, \mathrm will only work well on short items. Spaces are still not active and accented characters do not work^{4}.

4The AMS-L^{A}TEX package makes the \textrm command work with size changing.

3.7 Describing Variables  39

2nd 2^{nd} (3.10)

\begin{equation}
2^\textrm{nd} \quad
2^\mathrm{nd}
\end{equation}

Sometimes you need to tell L^{A}TEX about the correct font size nevertheless. In math mode the fontsize is set with the four commands:

\displaystyle (123), \textstyle (123), \scriptstyle (123) and
\scriptscriptstyle (123).

Changing styles also affects the way limits are displayed.

corr(X, Y) = \frac{∑_{i=1}^{n}(x_{i} − \overline{x})(y_{i} − \overline{y})}{[∑_{i=1}^{n}(x_{i} − \overline{x})^{2} ∑_{i=1}^{n}(y_{i} − \overline{y})^{2}]^{1/2}}

\begin{displaymath}
\mathop{\mathrm{corr}}(X,Y)=
 \frac{\displaystyle
   \sum_{i=1}^n(x_i-\overline x)
   (y_i-\overline y)}
  {\displaystyle\biggl[
 \sum_{i=1}^n(x_i-\overline x)^2
\sum_{i=1}^n(y_i-\overline y)^2
\biggr]^{1/2}}
\end{displaymath}

This is one of those examples in which we need smaller brackets than the standard \left[ \right] provides.

3.7 Describing Variables

For some formulae you might want to add a section after the formula where you describe the variables used in the expression. The following example should help you to get this done.

a^{2} + b^{2} = c^{2}

Where: a, b – are adjunct to the right angle of a right-angled triangle.

c – is the hypotenuse of the triangle.

\begin{displaymath}
a^2+b^2=c^2
\end{displaymath}
{\settowidth{\parindent}
   {Where:\ }
\makebox[0pt][r]
 {Where:\ }$a$, $b$ -- are
adjunct to the right angle
of a right-angled triangle.
$c$ -- is the hypotenuse of
the triangle.}

If you often need to typeset material like this, now the time is right to try out the \newenvironment command. Use it to create a specialised environment for describing variables. Check the description at the end of the previous chapter.

3.8 Theorems, Laws, …

When writing mathematical documents, you probably need a way to typeset “Lemmas”, “Definitions”, “Axioms” and similar structures. L^{A}TEX supports this with the command

\newtheorem{name}[counter]{text}[section]

The name argument is a short keyword used to identify the “theorem”. With the text argument you define the actual name of the “theorem” which will be printed in the final document.

The arguments in square brackets are optional. They are both used to specify the numbering used on the “theorem”. With the counter argument you can specify the name of a previously declared “theorem”. The new “theorem” will then be numbered in the same sequence. The section argument allows you to specify the sectional unit within which you want your “theorem” to be numbered.

After executing the \newtheorem command in the preamble of your document, you can use the following command within the document.

     \begin{name}[text]

This is my interesting theorem

     \end{name}

This should be enough theory. The following examples will hopefully remove the final remains of doubt, and make it clear that the \newtheorem environment is way too complex to understand.

Law 1 Don’t hide in the witness box

Jury 2 (The Twelve) It could be you! So beware and see law 1

Law 3 No, No, No

                                  % definitions for the document
                                  % preamble
                                  \newtheorem{law}{Law}
                                  \newtheorem{jury}[law]{Jury}
                                  %in the document
                                  \begin{law} \label{law:box}
                                  Don't hide in the witness box
                                  \end{law}
                                  \begin{jury}[The Twelve]
                                  It could be you! So beware and
                                  see law \ref{law:box}\end{jury}
                                  \begin{law}No, No, No\end{law}

The “Jury” theorem uses the same counter as the “Law” theorem. Therefore, it gets a number which is in sequence with the other “Laws”. The argument in square brackets is used to specify a title or something similar for the theorem.

Murphy 3.8.1 What can go wrong, will go wrong.

                                  \newtheorem{mur}{Murphy}[section]
                                  \begin{mur} What can go wrong,
                                  will go wrong. \end{mur}

The “Murphy” theorem gets a number which is linked to the number of the current section. You could also use another unit, like chapter or subsection for example.

3.9 Bold symbols

It is quite difficult to get bold symbols in L^{A}TEX; this is probably intentional as amateur typesetters tend to overuse them. The font change command \mathbf gives bold letters, but these are roman (upright) whereas mathematical symbols are normally italic. There is a \boldmath command, but this can only be used outside mathematics mode. It works for symbols too.

μ, M  M  μ, M

                                  \begin{displaymath}
                                  \mu, M \qquad \mathbf{M} \qquad
                                  \mbox{\boldmath $\mu, M$}
                                  \end{displaymath}

Notice that the comma is bold too, which may not be what is required.

The package amsbsy (included by amsmath) makes this much easier. It includes a \boldsymbol command, and “poor man’s bold” \pmb which works by over-striking for systems without the fonts required for bold symbols.

μ, M  μ, M  μ μ μ,M M M

                                  \begin{displaymath}
                                  \mu, M \qquad
                                  \boldsymbol{\mu}, \boldsymbol{M}
                                  \qquad \pmb{\mu}, \pmb{M}
                                  \end{displaymath}

3.10 List of Mathematical Symbols

In the following tables you find all the symbols normally accessible from math mode.

To use the symbols listed in Tables 3.12–3.16^{5}, the package amssymb must be loaded in the preamble of the document and the AMS math fonts must be installed on the system. If the AMS package and fonts are not installed, on your system, have a look at

CTAN:/tex-archive/macros/latex/packages/amslatex

Table 3.1: Math Mode Accents

 â \hat{a}     ǎ \check{a}   ã \tilde{a}     á \acute{a}
 à \grave{a}   ȧ \dot{a}     ä \ddot{a}      ă \breve{a}
 ā \bar{a}     a⃗ \vec{a}     Â \widehat{A}   Ã \widetilde{A}

Table 3.2: Lowercase Greek Letters

   α \alpha  θ \theta  o o  υ \upsilon
   β \beta  ϑ \vartheta π \pi  ϕ \phi
   γ \gamma  ι \iota  ϖ \varpi  φ \varphi
   δ \delta  κ \kappa  ρ \rho  χ \chi
   ϵ \epsilon  λ \lambda  ϱ \varrho  ψ \psi
   ε \varepsilon μ \mu  σ \sigma  ω \omega
   ζ \zeta  ν \nu  ς \varsigma
   η \eta  ξ \xi  τ \tau

Table 3.3: Uppercase Greek Letters

        Γ \Gamma Λ \Lambda Σ \Sigma  Ψ \Psi
        Δ \Delta Ξ \Xi  Υ \Upsilon Ω \Omega
        Θ \Theta Π \Pi  Φ \Phi

5These tables were derived from symbols.tex by David Carlisle and subsequently changed extensively as suggested by Josef Tkadlec

Table 3.4: Binary Relations

You can produce corresponding negations by adding a \not command as prefix to the following symbols.

< <  > >  = =
≤ \leq or \le  ≥ \geq or \ge  ≡ \equiv
≪ \ll  ≫ \gg           ≐ \doteq
≺ \prec  ≻ \succ  ∼ \sim
⪯ \preceq  ⪰ \succeq  ≃ \simeq
⊂ \subset  ⊃ \supset  ≈ \approx
⊆ \subseteq  ⊇ \supseteq     ≅ \cong
⊏ \sqsubset ^{a} ⊐ \sqsupset ^{a} ⋈ \Join ^{a}
⊑ \sqsubseteq ⊒ \sqsupseteq ▷◁ \bowtie
∈ \in  ∋ \ni , \owns  ∝ \propto
⊢ \vdash  ⊣ \dashv  ⊧ \models
| \mid  ∥ \parallel  ⊥ \perp
⌣ \smile  ⌢ \frown  ≍ \asymp
: :             ∉ \notin  ≠ \neq or \ne

aUse the latexsym package to access this symbol

Table 3.5: Binary Operators

+ +  − -
± \pm  ∓ \mp  ◁ \triangleleft
· \cdot  ÷ \div  ▷ \triangleright
× \times  \ \setminus  ⋆ \star
∪ \cup  ∩ \cap  ∗ \ast
⊔ \sqcup  ⊓ \sqcap  ∘ \circ
∨ \vee , \lor  ∧ \wedge , \land  • \bullet
⊕ \oplus  ⊖ \ominus  ⋄ \diamond
⊙ \odot  ⊘ \oslash  ⊎ \uplus
⊗ \otimes  ◯ \bigcirc  ⨿ \amalg
△ \bigtriangleup ▽ \bigtriangledown † \dagger
⊲ \lhd ^{a}  ⊳ \rhd ^{a}  ‡ \ddagger
⊴ \unlhd ^{a}  ⊵ \unrhd ^{a}  ≀ \wr

44  Typesetting Mathematical Formulae

Table 3.6: BIG Operators

     ∑ \sum  ⋃ \bigcup  ⋁ \bigvee  ⨁ \bigoplus
     ∏ \prod  ⋂ \bigcap  ⋀ \bigwedge ⨂ \bigotimes
     ∐ \coprod ⨆ \bigsqcup  ⨀ \bigodot
     ∫ \int  ∮ \oint  ⨄ \biguplus

Table 3.7: Arrows

← \leftarrow or \gets ⟵ \longleftarrow  ↑ \uparrow
→ \rightarrow or \to  ⟶ \longrightarrow  ↓ \downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
⇒ \Rightarrow  ⟹ \Longrightarrow  ⇓ \Downarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↩ \hookleftarrow  ↪ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀ \rightharpoonup  ↙ \swarrow
↽ \leftharpoondown  ⇁ \rightharpoondown  ↖ \nwarrow
⇌ \rightleftharpoons ⟺ \iff (bigger spaces)  ⇝ \leadsto ^{a}

aUse the latexsym package to access this symbol

Table 3.8: Delimiters

 ( (  ) )  ↑ \uparrow  ⇑ \Uparrow
 [ [ or \lbrack ] ] or \rbrack ↓ \downarrow  ⇓ \Downarrow
 { \{ or \lbrace } \} or \rbrace ↕ \updownarrow ⇕ \Updownarrow
 ⟨ \langle  ⟩ \rangle  | | or \vert  ∥ \| or \Vert
 ⌊ \lfloor  ⌋ \rfloor  ⌈ \lceil  ⌉ \rceil

/ /  \ \backslash  . (dual. empty)

Table 3.9: Large Delimiters

   { \lgroup     } \rgroup     { \lmoustache  } \rmoustache
    | \arrowvert ∥ \Arrowvert  | \bracevert

3.10 List of Mathematical Symbols  45

Table 3.10: Miscellaneous Symbols

. . . \dots  · · · \cdots       ⋮ \vdots  ... \ddots
ℏ \hbar  ı \imath  ȷ \jmath  ℓ \ell
ℜ \Re  ℑ \Im  ℵ \aleph  ℘ \wp
∀ \forall  ∃ \exists  ℧ \mho ^{a}  ∂ \partial
′ '  ′ \prime  ∅ \emptyset ∞ \infty
∇ \nabla  △ \triangle  □ \Box ^{a}  ◇ \Diamond ^{a}
⊥ \bot  ⊤ \top  ∠ \angle  √ \surd
♢ \diamondsuit ♡ \heartsuit ♣ \clubsuit ♠ \spadesuit
¬ \neg or \lnot ♭ \flat  ♮ \natural  ♯ \sharp

aUse the latexsym package to access this symbol

Table 3.11: Non-Mathematical Symbols

These symbols can also be used in text mode.

                † \dag  § \S   \begin{matrix}c \\ ◯ \end{matrix} \copyright
                ‡ \ddag ¶ \P £ \pounds

Table 3.12: AMS Delimiters

    ⌜ \ulcorner ⌝ \urcorner ⌞ \llcorner ⌟ \lrcorner

Table 3.13: AMS Greek and Hebrew

 𝟋 \digamma κ \varkappa ℶ \beth ℸ \daleth ℷ \gimel

Table 3.14: AMS Binary Relations

⋖ \lessdot  ⋗ \gtrdot  ≑ \doteqdot or \Doteq
⩽ \leqslant  ⩾ \geqslant  ≓ \risingdotseq
⪕ \eqslantless  ⪖ \eqslantgtr  ≒ \fallingdotseq
≦ \leqq  ≧ \geqq  ≖ \eqcirc
≪ \lll or \llless  ≫ \ggg or \gggtr  ≗ \circeq
≲ \lesssim  ≳ \gtrsim  ≜ \triangleq
⪅ \lessapprox  ⪆ \gtrapprox  ≏ \bumpeq
≶ \lessgtr  ≷ \gtrless  ≎ \Bumpeq
⋚ \lesseqgtr  ⋛ \gtreqless  ∼ \thicksim
⪋ \lesseqqgtr  ⪌ \gtreqqless  ≈ \thickapprox
≼ \preccurlyeq  ≽ \succcurlyeq  ≊ \approxeq
⋞ \curlyeqprec  ⋟ \curlyeqsucc  ∽ \backsim
≾ \precsim  ≿ \succsim  ⋍ \backsimeq
≾ \precapprox  ≿ \succapprox  ⊨ \vDash
⫅ \subseteqq  ⫆ \supseteqq  ⊩ \Vdash
⋐ \Subset  ⋑ \Supset  ⊪ \Vvdash
⊏ \sqsubset  ⊐ \sqsupset  ϶ \backepsilon
∴ \therefore  ∵ \because  ∝ \varpropto
 ∣ \shortmid  ∥ \shortparallel  ≬ \between
⌣ \smallsmile  ⌢ \smallfrown  ⋔ \pitchfork
◁ \vartriangleleft ▷ \vartriangleright ◀ \blacktriangleleft
⊴ \trianglelefteq  ⊵ \trianglerighteq  ▶ \blacktriangleright

Table 3.15: AMS Arrows

←−− \dashleftarrow  −−→ \dashrightarrow  ⊸ \multimap
⇇ \leftleftarrows  ⇉ \rightrightarrows  ⇈ \upuparrows
⇆ \leftrightarrows  ⇄ \rightleftarrows  ⇊ \downdownarrows
⇚ \Lleftarrow  ⇛ \Rrightarrow  ↿ \upharpoonleft
↞ \twoheadleftarrow  ↠ \twoheadrightarrow  ↾ \upharpoonright
↢ \leftarrowtail  ↣ \rightarrowtail  ⇃ \downharpoonleft
⇋ \leftrightharpoons ⇌ \rightleftharpoons  ⇂ \downharpoonright
↰ \Lsh  ↱ \Rsh  ⇝ \rightsquigarrow
↫ \looparrowleft  ↬ \looparrowright  ↭ \leftrightsquigarrow
↶ \curvearrowleft  ↷ \curvearrowright
⟲ \circlearrowleft  ⟳ \circlearrowright

Table 3.16: AMS Negated Binary Relations and Arrows

≮ \nless  ≯ \ngtr  ⫋ \varsubsetneqq
⪇ \lneq  ⪈ \gneq  ⫌ \varsupsetneqq
≰ \nleq  ≱ \ngeq  ⫅̸ \nsubseteqq
⩽̸ \nleqslant  ⩾̸ \ngeqslant  ⫆̸ \nsupseteqq
≨ \lneqq  ≩ \gneqq  ∤ \nmid
≨ \lvertneqq  ≩ \gvertneqq  ∦ \nparallel
≦̸ \nleqq  ≧̸ \ngeqq  ∤ \nshortmid
⋦ \lnsim  ⋧ \gnsim  ∦ \nshortparallel
⪉ \lnapprox  ⪊ \gnapprox  ≁ \nsim
⊀ \nprec  ⊁ \nsucc  ≇ \ncong
⪯̸ \npreceq  ⪰̸ \nsucceq  ⊬ \nvdash
⪵ \precneqq  ⪶ \succneqq  ⊭ \nvDash
⋨ \precnsim  ⋩ \succnsim  ⊮ \nVdash
⪹ \precnapprox  ⪺ \succnapprox  ⊯ \nVDash
⊊ \subsetneq  ⊋ \supsetneq  ⋪ \ntriangleleft
⊊ \varsubsetneq ⊋ \varsupsetneq ⋫ \ntriangleright
⊈ \nsubseteq  ⊉ \nsupseteq  ⋬ \ntrianglelefteq
⫋ \subsetneqq  ⫌ \supsetneqq  ⋭ \ntrianglerighteq
↚ \nleftarrow  ↛ \nrightarrow  ↮ \nleftrightarrow
⇍ \nLeftarrow  ⇏ \nRightarrow  ⇎ \nLeftrightarrow

Table 3.17: AMS Binary Operators

∔ \dotplus  ▪ \centerdot  ⊺ \intercal
⋉ \ltimes  ⋊ \rtimes  ⋇ \divideontimes
⋓ \Cup or \doublecup ⋒ \Cap or \doublecap ∖ \smallsetminus
⊻ \veebar  ⊼ \barwedge  ⩞ \doublebarwedge
⊞ \boxplus  ⊟ \boxminus  ⊝ \circleddash
⊠ \boxtimes  ⊡ \boxdot  ⊚ \circledcirc
⋋ \leftthreetimes  ⋌ \rightthreetimes ⊛ \circledast
⋎ \curlyvee  ⋏ \curlywedge

Table 3.18: AMS Miscellaneous

ℏ \hbar  ℏ \hslash  𝕜 \Bbbk
□ \square  ■ \blacksquare  Ⓢ \circledS
△ \vartriangle  ▲ \blacktriangle  ∁ \complement
▽ \triangledown ▼ \blacktriangledown ⅁ \Game
♢ \lozenge  ♦ \blacklozenge  ⋆ \bigstar
∠ \angle  ∡ \measuredangle  ∢ \sphericalangle
⧸ \diagup  ⧹ \diagdown  ‵ \backprime
∄ \nexists  Ⅎ \Finv  ∅ \varnothing
ð \eth  ℧ \mho

Table 3.19: Math Alphabets

Example Command  Required package

ABCdef \mathrm{ABCdef}
ABCdef \mathit{ABCdef}
ABCdef \mathnormal{ABCdef}
ABC  \mathcal{ABC}

𝒜ℬ𝒞  \mathcal{ABC}  eucal with option: mathcal
or  \mathscr{ABC}  eucal with option: mathscr

𝔄𝔅ℭ𝔡𝔢𝔣 \mathfrak{ABCdef}  eufrak

𝔸𝔹ℂ  \mathbb{ABC}  amsfonts or amssymb

Specialities

Don’t read this chapter! — Or, at least, if you feel confident enough, you can now start writing your documents in L^{A}TEX. The purpose of this chapter is to add some ‘spice’ to your L^{A}TEX knowledge. A much more complete description of specialities and enhancements possible with L^{A}TEX can be found in the L^{A}TEX Manual [1] and The L^{A}TEX Companion [3].

4.1 Fonts and Sizes

L^{A}TEX chooses the appropriate font and font size based on the logical structure of the document (sections, footnotes, …). In some cases one might like to change fonts and sizes by hand. To do this you can use the commands listed in Tables 4.1 and 4.2. The actual size of each font is a design issue and depends on the document class and its options.

The small and bold Romans ruled

all of great big Italy.

                                  {\small The small and
                                  \textbf{bold} Romans ruled}
                                  {\Large all of great big
                                  \textit{Italy}.}

One important feature of L^{A}TEX 2ε is, that the font attributes are independent. This means, that you can change issue size or even font changing commands and still keep the bold or slant attribute set earlier. While this may be evident to someone learning L^{A}TEX from scratch, it is not all that clear the someone used to L^{A}TEX 2.09.

In math mode you can use the font changing commands to temporarily exit math mode and enter some normal text. If you want to switch to another font for math typesetting there exists another special set of commands. Refer to Table 4.3.

In connection with the font size commands, curly braces play a significant role. They are used to to build groups. Groups limit the scope of most L^{A}TEX commands.

Table 4.1: Fonts

\textrm{...} roman  \textsf{...} sans serif
\texttt{...}  typewriter
\textmd{...} medium  \textbf{...} bold face
\textup{...} upright  \textit{...} italic
\textsl{...} slanted  \textsc{...} small caps
  \emph{...} emphasised  \textnormal{...} document font

Table 4.2: Font sizes

\tiny  tiny font
\scriptsize  very small font
\footnotesize quite small font
\small  small font

\normalsize  normal font

\large  large font

\Large larger font
\LARGE very large font

                 \huge huge

\Huge largest

Table 4.3: Math fonts

Command  Example  Output

\mathcal{...}     $\mathcal{B}=c$  B = c
\mathrm{...}      $\mathrm{K}_2$  K_{2}
\mathbf{...}      $\sum x=\mathbf{v}$  ∑ x = v
\mathsf{...}      $\mathsf{G\times R}$  G × R
\mathtt{...}      $\mathtt{L}(b,c)$                 L(b, c)
\mathnormal{...}  $\mathnormal{R_{19}}\neq R_{19}$ R_{19} ≠ R_{19}
\mathit{...}      $\mathit{ffi}\neq ffi$  ffi ≠ ffi

\underline{4.2 Spacing 51}

He likes large and small letters.

                                  He likes {\LARGE large and
                                  {\small small} letters}.

The font size commands also change the line spacing, but only if the paragraph ends within the scope of the font size command. The closing curly brace } should therefore not come too early. Note the position of the \par command in the next two examples.

Don’t read this! It is not true. You can believe me!

                                  {\Large Don't read this! It is not
                                  true. You can believe me!\par}

This is not true. But remember I am a liar.

                                  {\Large This is not true.
                                  But remember I am a liar.}\par

To conclude this journey into the land of fonts and font sizes, here is a little word of advice^{1}:

\underline{Remember!} The MO RE fonts you use in a document the more readable and beautiful it becomes.

4.2 Spacing

4.2.1 Line Spacing

If you want to use larger inter line spacing in a document, you can change its value by putting the

 \linespread{factor}

command into the preamble of your document. Use \linespread{1.3} for “one and a half” line spacing, and \linespread{1.6} for “double” line spacing. Normally the lines are not spread, therefore the default line spread factor is 1.

4.2.2 Paragraph Formating

In L^{A}TEX, there are two parameters influencing paragraph layout. By placing a definition like

     \setlength{\parindent}{0pt}
     \setlength{\parskip}{1ex plus 0.5ex minus 0.2ex}

1Attention: This is a bit of satire. I hope you realise that!

in the preamble^{2} of the input file the appearance of paragraphs can be changed. These two lines increase the space between two paragraphs while setting the paragraph indent to zero. In continental Europe, paragraphs are often separated by some space and not indented. But beware, this has also its effect on the table of contents. Its lines get spaced more loosely now as well.

If you want to indent a paragraph which is not indented, you can use

 \indent

at the beginning of the paragraph^{3}. Obviously this can only work, when \parindent is not set to zero.

To create a non-indented paragraph you can use

 \noindent

as the first command of the paragraph. This might come in handy, when you start a document with body text and not with a sectioning command.

4.2.3 Horizontal Space

L^{A}TEX determines the spaces between words and sentences automatically. To add horizontal space, use:

 \hspace{length}

If such a space should be kept, even if it falls at the end or the start of a line, use \hspace* instead of \hspace. The length in the simplest case just is a number plus a unit. The most important units are listed in Table 4.4.

This is a space of 1.5 cm.

                                  This\hspace{1.5cm}is a space
                                  of 1.5 cm.

The command

 \stretch{n}

generates a special rubber space. It stretches, until all the remaining space on a line is filled up. If two \hspace{\stretch{n}} commands are issued on the same line, they grow according to the stretch factor.

2Between the \documentclass and the \begin{document} commands

3To indent the first paragraph after each section head, use the indentfirst package in the ‘tools’ bundle

Table 4.4: TEX Units

mm millimetre ≈ 1/25 inch
cm centimetre = 10 mm
in inch ≈ 25.4 mm
pt point ≈ 1/72 inch ≈ \frac{1}{3} mm
em approx width of an m in the current font
ex approx height of an x in the current font

x  x  x

                                  x\hspace{\stretch{1}}
                                  x\hspace{\stretch{3}}x

4.2.4 Vertical Space

The space between paragraphs, sections, subsections, … is determined automatically by L^{A}TEX. If necessary, additional vertical space between two paragraphs can be added with the command:

 \vspace{length}

This command should normally be used between two empty lines. If the space should be preserved at the top, or at the bottom of a page, use the starred version of the command \vspace* instead of \vspace.

The \stretch command in connection with \pagebreak can be used to typeset text on the last line of a page, or to centre text vertically on a page.

     Some text \ldots
     \vspace{\stretch{1}}
     This goes onto the last line of the page.\pagebreak

Additional space between two lines of the same paragraph or within a table is specified with the

\\[length]

command.

4.3 Page Layout

L^{A}TEX 2ε allows you to specify the paper size in the \documentclass command. It then automatically picks the right text margins. But some times,

54  Specialities

[figure: Header; Body; Footer; Margin Notes; 8; 7; 1; 3; 10; 9; 11; 2; 4; 5; 6]

1   one inch + \hoffset       2   one inch + \voffset
3   \oddsidemargin = 22pt     4   \topmargin = 22pt
5   \headheight = 13pt        6   \headsep = 19pt
7   \textheight = 595pt       8   \textwidth = 360pt
9   \marginparsep = 7pt      10   \marginparwidth = 106pt
11   \footskip = 27pt              \marginparpush = 5pt (not shown)
    \hoffset = 0pt                \voffset = 0pt
    \paperwidth = 597pt           \paperheight = 845pt

Figure 4.1: Page Layout Parameters you may not be happy with the predefined values. Naturally, you can change them. Figure 4.1 shows all the parameters which can be changed. The figure was produced with the layout package from the tools bundle^{4}.

L^{A}TEX provides two commands to change these parameters. They are usually used in the document preamble.

The first command assigns a fixed value to any of the parameters:

\setlength{parameter}{length}

The second command, adds a length to any of the parameters.

 \addtolength{parameter}{length}

This second command is actually more useful than the \setlength command, because you can now work relative to the existing settings. To add one centimetre to the overall text width, I would put the following commands into the document preamble:

     \addtolength{\hoffset}{-0.5cm}
     \addtolength{\textwidth}{1cm}

4.4 Bibliography

With the thebibliography environment you can produce a bibliography. Each entry starts with

 \bibitem{marker}

The marker is then used to cite the book within the document.

 \cite{marker}

The numbering of the entries is generated automatically. The parameter after the \begin{thebibliography} command sets the maximum width of these numbers.

4CTAN:/tex-archive/macros/latex/packages/tools

Partl [1] has proposed, that …

Bibliography

[1] H. Partl: German TEX, TUGboat Vol. 9, No. 1 (’88)

                                  Partl~\cite{pa} has
                                  proposed, that \ldots
                                  \begin{thebibliography}{99}
                                  \bibitem{pa} H.~Partl:
                                  \emph{German \TeX},
                                  TUGboat Vol.~9, No.~1 ('88)
                                  \end{thebibliography}

4.5 Indexing

A very useful feature of many books is their index. With L^{A}TEX and the support program makeindex^{5} indexes can be generated quite easily. In this introduction, only the basic index generation commands will be explained. For a more in depth view please refer to The L^{A}TEX Companion [3].

To enable the indexing feature of L^{A}TEX the makeidx package must be loaded in the preamble with:

 \usepackage{makeidx}

and the special indexing commands must be enabled by putting the

 \makeindex

command into the input files preamble.

The contents of the index is specified with

 \index{key}

commands. Where key is the index entry. You enter the index commands at the points in the text where you want the final index entries to point to. Table 4.5 explains the syntax of the key argument with several examples.

When the input file is processed with L^{A}TEX, each \index command writes an appropriate index entry together with the current page number to a special file. The file has the same name as the L^{A}TEX input file, but a

5On systems not supporting filenames longer than 8 characters, the program may be called makeidx.

Table 4.5: Index Key Syntax Examples

Example  Index Entry Comment
\index{hello}  hello, 1  Plain entry
\index{hello!Peter}  Peter, 3  Subentry under ‘hello’

  \index{Sam@\textsl{Sam}} Sam, 2  Formated entry
  \index{Lin@\textbf{Lin}} Lin, 7  Same as above

\index{Jenny|textbf}  Jenny, 3  Formated page number

  \index{Joe|textit}  Joe, 5  Same as above

different extension (.ind). This .ind file can then be processed with the

makeindex program.
 makeindex filename

The makeindex program generates a sorted index with the same base file name, but this time with the extension .idx. If now the L^{A}TEX input file is processed again, this sorted index gets included into the document at the point where L^{A}TEX finds

 \printindex

The showidx package which comes with L^{A}TEX 2ε prints out all index entries in the left margin of the text. This is quite useful for proofreading a document and verifying the index.

4.6 Including EPS Graphics

With the figure and the table environment L^{A}TEX provides the basic facilities to work with floating bodies such as images or graphics.

There are also several possibilities to generate the actual graphics with basic L^{A}TEX or a L^{A}TEX extension package. Unfortunately, most users find them quite difficult to understand. Therefore this will not be explained any further in this manual. For more information on that subject please refer to The L^{A}TEX Companion [3] and the L^{A}TEX Manual [1].

A much easier way to get graphics into a document, is to generate them with a specialised software package^{6} and then include the finished graphics into the document. Here again, L^{A}TEX packages offer many ways to do that. In this introduction, only the use of Encapsulated PostScript (EPS) graphics will be discussed, because it is quite easy to do and widely used. In order to use pictures in the EPS format, you must have a PostScript printer^{7} available for output.

6Such as XFig, CorelDraw!, Freehand, Gnuplot, …

A good set of commands for inclusion of graphics is provided in the graphicx package by D. P. Carlisle. It is part of a whole family of packages called the “graphics” bundle^{8}.

Assuming you are working on a system with a PostScript printer available for output and with the graphicx package installed, you can use the following step by step guide to include a picture into your document:

1. Export the picture from your graphics program in EPS format.

2. Load the graphicx package in the preamble of the input file with

 \usepackage[driver]{graphicx}

driver is the name of your “dvi to postscript” converter^{9}. This information is required by the package because the actual graphics inclusion is done by the printer driver. Knowing the driver, the graphicx package inserts the correct commands into the .dvi file for the printer driver to include the desired EPS graphics.

3. Use the command

 \includegraphics[key=value, . . . ]{file}

to include file into your document. The optional parameter accepts a comma separated list of keys and associated values. The keys can be used to alter the width, height and rotation of the included graphic. Table 4.6 lists the most important keys.

Table 4.6: Key Names for graphicx Package

width scale graphic to the specified width
height scale graphic to the specified height
angle rotate graphics clockwise

The following example code will hopefully make things clear:

7Another possibility to output PostScript is the GhostScript program available from CTAN:/tex-archive/support/ghostscript

8CTAN:/tex-archive/macros/latex/packages/graphics 9The most widely used program is called dvips.

\begin{figure}
\begin{center}
\includegraphics[angle=90, width=10cm]{test.eps}
\end{center}
\end{figure}

This includes the graphic stored in the file test.eps. The graphic is first rotated by 90 degrees and then scaled to the final width of 10 cm. The aspect ratio is 1.0 because no special height is specified.

For more information please refer to [8].

Bibliography

[1] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1994, ISBN 0-201-52983-1.

[2] Donald E. Knuth. The TEXbook, Volume A of Computers and Typesetting, Addison-Wesley Publishing Company (1984), ISBN 0-201-13448-9.

[3] Michel Goossens, Frank Mittelbach and Alexander Samarin. The L^{A}TEX Companion. Addison-Wesley, Reading, Massachusetts, 1994, ISBN 0-201-54199-8.

[4] Each L^{A}TEX installation should provide a so-called L^{A}TEX Local Guide, which explains the things which are special to the local system. It should be contained in file called local.tex. Unfortunately some lazy sysops do not provide such a document. In this case, go and ask you local L^{A}TEX guru for help.

[5] L^{A}TEX3 Project Team. L^{A}TEX 2ε for authors. Comes with the L^{A}TEX 2ε distribution as usrguide.tex.

[6] L^{A}TEX3 Project Team. L^{A}TEX 2ε for Class and Package writers. Comes with the L^{A}TEX 2ε distribution as clsguide.tex.

[7] L^{A}TEX3 Project Team. L^{A}TEX 2ε Font selection. Comes with the L^{A}TEX 2ε distribution as fntguide.tex.

[8] D. P. Carlisle. Packages in the ‘graphics’ bundle. Comes with the ‘graphics’ bundle as grfguide.tex, available form the same source your L^{A}TEX distribution came from.

Index

\!, 36 ", 15 $, 31 ', 34 \(, 31 \), 31 \,, 32, 36 -, 15 −, 15 \-, 14 –, 15 —, 15 ., space after, 18 …, 16 \@, 18 \[, 31 \\, 13, 22–24, 53

\\*, 13

\], 31 ^, 33 _, 33 |, 24 ~, 18

A4 paper, 8 A5 paper, 8 accent, 16 acute, 17

\addtolength, 55

advantages of L^{A}TEX, 3 æ, 17 amsbsy, 41 amsfonts, 33 amsmath, 36, 38, 41 amssymb, 33, 42

\and, 19

                                 \appendix, 18, 19
                                  array, 37

arrow symbols, 34 article class, 8

                                 \atop, 35
                                 \author, 19
                                 \authors, 19

B5 paper, 8 babel, 17

                                 \backmatter, 20

backslash, 5 base font size, 8

                                 \begin, 21
                                 \bibitem, 55

bibliography, 55

                                 \Big, 36
                                 \big, 36
                                 \Bigg, 36
                                 \bigg, 36

blackboard bold, 33

                                 \bmod, 35

bold face, 50

bold symbols, 33, 41

                                 \boldmath, 41
                                 \boldsymbol, 41

book class, 8 braces, 35

                                 \caption, 27
                                 \cdots, 36
                                  center, 22
                                 \chapter, 18
                                 \choose, 35
                                 \cite, 55
                                 \cleardoublepage, 28

\underline{INDEX 63}

\clearpage, 28

coloured text, 9 comma, 16 commands, 5

\!, 36 \(, 31 \), 31 \,, 32, 36 \-, 14 \@, 18 \[, 31 \\, 13, 22–24, 53

\\*, 13

\], 31

\addtolength, 55
\and, 19
\appendix, 18, 19
\atop, 35
\author, 19
\authors, 19
\backmatter, 20
\begin, 21
\bibitem, 55
\Big, 36
\big, 36
\Bigg, 36
\bigg, 36
\bmod, 35
\boldmath, 41
\boldsymbol, 41
\caption, 27
\cdots, 36
\chapter, 18
\choose, 35
\cite, 55
\cleardoublepage, 28
\clearpage, 28
\date, 19
\ddots, 36
\displaystyle, 39
\documentclass, 7

\emph, 21, 50

\end, 21
\footnote, 20

\footnotesize, 50
\frac, 35
\frenchspacing, 18
\frontmatter, 20
\fussy, 14
\hline, 24
\hspace, 52
\Huge, 50
\huge, 50
\hyphenation, 14
\include, 11, 12
\includegraphics, 58
\includeonly, 12
\indent, 52
\index, 56
\input, 12
\int, 35
\item, 21
\label, 20, 32
\LARGE, 50
\Large, 50
\large, 50
\ldots, 16, 36
\left, 36
\linebreak, 14
\linespread, 51
\listoffigures, 27
\listoftables, 27
\mainmatter, 20
\makeindex, 56
\maketitle, 19
\mathbb, 33
\mathbf, 50
\mathcal, 50
\mathit, 50
\mathnormal, 50
\mathrm, 38, 50
\mathsf, 50
\mathtt, 50

\mbox, 15, 16

\multicolumn, 25
\newcommand, 28, 29
\newenvironment, 29, 30
\newline, 13

\underline{64 INDEX}

\newpage, 13
\newtheorem, 40
\noindent, 52
\nolinebreak, 14
\nonumber, 38
\nopagebreak, 14
\normalsize, 50
\overbrace, 34
\overleftarrow, 34
\overline, 34
\overrightarrow, 34
\pagebreak, 14
\pageref, 20
\pagestyle, 11
\paragraph, 18
\parindent, 51
\parskip, 51

\part, 18, 19

\pmb, 41
\pmod, 35
\printindex, 57
\providecommand, 29
\qquad, 32, 36

\quad, 32, 36 \ref, 20, 32

\renewcommand, 29
\renewenvironment, 30
\right, 36, 37
\right., 36
\scriptscriptstyle, 39
\scriptsize, 50
\scriptstyle, 39
\section, 18
\setlength, 51, 55
\sloppy, 14
\small, 50
\sqrt, 34
\stretch, 52
\subparagraph, 18
\subsection, 18
\subsubsection, 18
\sum, 35
\tableofcontents, 19
\textbf, 50

                                  \textit, 50
                                  \textmd, 50
                                  \textnormal, 50
                                  \textrm, 38, 50
                                  \textsc, 50
                                  \textsf, 50
                                  \textsl, 50
                                  \textstyle, 39
                                  \texttt, 50
                                  \textup, 50
                                  \thispagestyle, 11
                                  \tiny, 50
                                  \title, 19
                                  \underbrace, 34
                                  \underline, 34
                                  \usepackage, 9, 17
                                  \vdots, 36
                                  \vec, 34

\verb, 23, 24

                                  \vspace, 53
                                  \widehat, 34
                                  \widetilde, 34

comments, 5

cross references, 20 curly braces, 5, 49

dash, 15

                              \date, 19

dcolumn, 25

                              \ddots, 36

decimal alignment, 25 delimiters, 35 describing variables, 40

                               description, 21

diagonal dots, 36 dimensions, 52

                               displaymath, 31
                              \displaystyle, 39

document font size, 8 document title, 8

                              \documentclass, 7

dotless ıand ȷ, 17

double line spacing, 51 double sided, 8 ellipsis, 16 em-dash, 15 \emph, 21, 50 emphasise, 21

empty, 11

en-dash, 15 Encapsulated PostScript, 57

\end, 21
enumerate, 21

environment, 21 environments

   array, 37
   center, 22
   description, 21
   displaymath, 31
   enumerate, 21
   eqnarray, 37
   equation, 32
   figure, 26, 27
   flushleft, 22
   flushright, 22
   itemize, 21
   math, 31
   quotation, 23
   quote, 22

table, 26, 27

   tabular, 24
   thebibliography, 55
   verbatim, 23
   verse, 23
eqnarray, 37
equation, 32

equation system, 37 executive paper, 8 exponent, 33 exscale, 36

figure, 26, 27

floating bodies, 26

flushleft, 22
flushright, 22

font, 49 font encoding, 10 font size, 49, 50 footer, 11

                                 \footnote, 20
                                 \footnotesize, 50

formulae, 31

                                 \frac, 35

fraction, 35

                                 \frenchspacing, 18
                                 \frontmatter, 20
                                 \fussy, 14

German, 17 GhostScript, 58 graphics, 9, 57 graphicx, 58 grave, 17 Greek letters, 33 grouping, 49

header, 11

textttheadings, 11

                                 \hline, 24

horizontal brace, 34 dots, 36 line, 34 space, 52

                                 \hspace, 52
                                 \Huge, 50
                                 \huge, 50

hyphen, 15

                                 \hyphenation, 14

hyphenation rules, 17

                                 \include, 11, 12
                                 \includegraphics, 58
                                 \includeonly, 12
                                 \indent, 52

indentfirst, 52 index, 56

                                 \index, 56
                                 \input, 12

input file, 6 inputenc, 17

                                 \int, 35

integral operator, 35 international, 17 italic, 50

\item, 21
itemize, 21

Knuth, Donald E., 1

\label, 20, 32

Lamport, Leslie, 1 language, 17

\LARGE, 50
\Large, 50
\large, 50

L^{A}TEX 2.09, 1 L^{A}TEX 2ε, 1 L^{A}TEX3, 1, 4 layout, 55

\ldots, 16, 36
\left, 36

left aligned, 22 legal paper, 8 letter paper, 8 ligature, 16 line spacing, 51 linebreak, 13

\linebreak, 14
\linespread, 51
\listoffigures, 27
\listoftables, 27

long equations, 38

\mainmatter, 20

makeidx, 56 makeidx package, 56

\makeindex, 56

makeindex program, 56

\maketitle, 19

margins, 53

math, 31

math font size, 38 math spacing, 36

\mathbb, 33
\mathbf, 50
\mathcal, 50

mathematical accents, 34 delimiter, 36 functions, 34 minus, 15 mathematics, 31

                                 \mathit, 50
                                 \mathnormal, 50
                                 \mathrm, 38, 50
                                 \mathsf, 50
                                 \mathtt, 50

\mbox, 15, 16 minus sign, 15 Mittelbach, Frank, 1 modulo function, 35

                                 \multicolumn, 25
                                 \newcommand, 28, 29
                                 \newenvironment, 29, 30
                                 \newline, 13
                                 \newpage, 13
                                 \newtheorem, 40
                                 \noindent, 52
                                 \nolinebreak, 14
                                 \nonumber, 38
                                 \nopagebreak, 14
                                 \normalsize, 50

œ, 17 option, 7 optional parameters, 5

                                 \overbrace, 34

overfull hbox, 14

                                 \overleftarrow, 34
                                 \overline, 34
                                 \overrightarrow, 34

package, 6, 7 packages

amsbsy, 41 amsfonts, 33 amsmath, 36, 38, 41 amssymb, 33, 42 babel, 17 dcolumn, 25 exscale, 36 graphicx, 58 indentfirst, 52 inputenc, 17 layout, 55 makeidx, 56 showidx, 57

page layout, 53 page style, 11

empty, 11
headings, 11
plain, 11
\pagebreak, 14
\pageref, 20
\pagestyle, 11

paper size, 8, 53

\paragraph, 18

parameter, 5

\parindent, 51
\parskip, 51

\part, 18, 19 period, 16 placement specifier, 26

plain, 11
\pmb, 41
\pmod, 35

PostScript, 58 preamble, 6 prime, 34

\printindex, 57
\providecommand, 29
\qquad, 32, 36

\quad, 32, 36

quotation, 23

quotation marks, 15

quote, 22

\ref, 20, 32

\renewcommand, 29
\renewenvironment, 30

report class, 8 reserved characters, 4

\right, 36, 37

right aligned, 22

\right., 36

roman, 50

sans serif, 50 Scandinavian letters, 17

\scriptscriptstyle, 39
\scriptsize, 50
\scriptstyle, 39
\section, 18
\setlength, 51, 55

showidx, 57 single sided, 8 slanted, 50 slide class, 8

\sloppy, 14
\small, 50

small caps, 50 space, 4 special character, 16

\sqrt, 34

square brackets, 5 square root, 34

\stretch, 52

structure, 6

\subparagraph, 18

subscript, 33

\subsection, 18
\subsubsection, 18
\sum, 35

sum operator, 35

table, 24 table, 26, 27 table of contents, 19

\tableofcontents, 19
 tabular, 24
\textbf, 50
\textit, 50
\textmd, 50
\textnormal, 50
\textrm, 38, 50
\textsc, 50
\textsf, 50
\textsl, 50

\textstyle, 39
\texttt, 50
\textup, 50
thebibliography, 55
\thispagestyle, 11

three dots, 36 tilde, 34 tilde ( ~), 18

\tiny, 50

title, 8, 19

\title, 19

two column, 8

umlaut, 17

\underbrace, 34

underfull hbox, 14

\underline, 34

units, 52, 53 upright, 50

\usepackage, 9, 17
\vdots, 36
\vec, 34

vectors, 34 \verb, 23, 24

verbatim, 23
verse, 23

vertical dots, 36 vertical space, 53

\vspace, 53

whitespace, 4 after commands, 5 at the start of a line, 4

\widehat, 34
\widetilde, 34

WYSIWYG, 2, 3

INDEX  69